### v2

# NdLinear: Don’t Flatten! Building Superior Neural Architectures by Preserving N-D Structure

Abstract
Many high-impact machine learning tasks involve multi-dimensional data such as images, volumetric medical scans, and multivariate time-series. Yet, most neural architectures flatten these inputs, discarding critical cross-dimension information. We introduce NdLinear, a novel linear transformation that circumvents this destructive flattening by operating directly on tensors. NdLinear applies transformations separately along each data dimension, thereby preserving the native data structure. Extensive experiments demonstrate NdLinear’s capacity to significantly enhance representational power, achieve dramatic parameter reductions (often by orders of magnitude), and maintain a favorable computational profile. For instance, when applied to Large Language Model finetuning, our NdLinear-LoRA delivers comparable or improved accuracy on reasoning tasks using up to 9×9$\times$9 × fewer trainable parameters than standard LoRA. These broad advantages of NdLinear are consistently validated across diverse neural architectures (CNNs, RNNs, Transformers, MLPs) and data domains, including vision, language, time-series, and tabular tasks. As a versatile, drop-in replacement for standard linear layers, NdLinear processes data in its original N-dimensional form, offering a foundational component for developing more efficient and powerful next-generation neural architectures.
footnotetext: NdLinear is licensed by Ensemble AI under the Apache 2.0 license. Our technical contributions and research findings are free for use by the open-source community.
1 Introduction
Effectively processing N-dimensional (N-D) tensors is critical in modern artificial intelligence, with applications spanning volumetric medical scans (Milletari et al., 2016), spatio-temporal video, multivariate time series (Liu et al., 2023), and outputs of complex simulations (Shanahan et al., 2023; Brown et al., 2020; Dosovitskiy et al., 2021; Radford et al., 2021). However, a core challenge remains: fully retaining and leveraging the rich, inherent structure of N-D data within neural architectures. While specialized layers like convolutions excel at capturing local spatial hierarchies (e.g., in 2D/3D images), they are not universally optimal, especially for comprehensively propagating information in higher-dimensional data (e.g., >3D) without excessive complexity. Fundamentally, many neural architectures (CNNs, Transformers, MLPs and RNNs) still rely on standard Linear (fully connected) layers, which expect 1D vector inputs. This common practice necessitates flattening multi-dimensional feature maps or tensors before these transformations, for instance, in network heads, channel-wise projections, or MLP blocks within attention mechanisms.
This ubiquitous flattening, even if localized, is often detrimental. It discards the data’s explicit N-D organization, disrupting critical spatial, temporal, or other relationships encoded across dimensions (Sokar et al., 2024). For example, valuable spatial hierarchies preserved by convolutions can be lost when feature vectors are decontextualized by flattening. In multivariate time series (e.g., dimensions for time, sensors, features), flattening intermingles these distinct axes, forcing models to rediscover temporal patterns or inter-sensor dependencies without structural priors. This loss of explicit structure forces networks to relearn intrinsic dependencies, potentially increasing parameter counts, data requirements, and computational load, while risking spurious correlations.
To address this, we introduce NdLinear, a novel linear layer that operates natively on N-D tensors. Unlike standard linear layers, NdLinear processes data in its original N-D form by applying distinct linear transformations sequentially along each tensor dimension. This dimension-wise approach inherently preserves the data’s N-D structure.
Preserving data’s native organization allows NdLinear to offer compelling advantages: It enhances representational power by maintaining structural integrity and facilitating a more natural information flow. It dramatically reduces parameter counts (often by orders of magnitude vs. flattened layers) and ensures efficient computation without sacrificing performance. These efficiencies lead to faster training/inference and lower memory usage. As a general-purpose layer, NdLinear is a versatile, drop-in replacement for standard linear layers in many architectures, benefiting diverse domains by enabling models to better leverage N-D data characteristics.
Our work makes the following primary contributions, substantiated by extensive empirical validation:
  • We articulate the critical limitations of flattening in conventional linear layers for N-D data.
  • We introduce and formulate NdLinear, a novel N-D linear layer using sequential, dimension-wise transformations to preserve data structure and achieve substantial parameter efficiency.
  • We empirically demonstrate through comprehensive evaluations that NdLinear significantly enhances model performance, or matches it with drastically fewer parameters, when replacing standard linear layers across diverse architectures (CNNs, RNNs, Transformers, MLPs) and data domains, including vision, language, time-series, and tabular data.
  • We analyze the significant parameter and computational efficiency gains originating from NdLinear’s dimension-wise transformation principle.
  • We establish NdLinear as a versatile and foundational building block for developing more efficient, structurally-informed, and powerful neural architectures.
The remainder of this paper is organized as follows. Section 2 details the NdLinear transformation. Section 3 analyzes its parameter efficiency, computational complexity, and representational capacity. Section 4 presents experimental results demonstrating NdLinear’s efficacy. Finally, Section 5 summarizes findings, discusses implications, and outlines future work.
2 NdLinear: N-dimensional Linear Transformation
This section formally introduces the NdLinear transformation. We first define the general problem of transforming tensor-structured data. Then, we detail the NdLinear layer, including its operational mechanics and algorithm. Finally, we situate NdLinear within the context of related methods.
2.1 Problem Formulation
This work addresses learnable linear transformations for N-dimensional (N-D) input tensors. Such an input, $X\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$, comprises a batch of $B$ samples, each with $N$ feature dimensions of sizes $D_{1},\ldots,D_{N}$ (e.g., height, width, channels for images; sequence length, embedding dimensions for text). The objective is to transform $X$ into an output tensor $Y^{\prime}\in\mathbb{R}^{B\times H_{1}\times\cdots\times H_{N}}$, where each $k$-th feature dimension is reshaped from its original size $D_{k}$ to a new size $H_{k}$. Our focus is specifically on the formulation of this N-D linear transformation.
A primary challenge is that conventional linear layers necessitate flattening $X$ into a 2D matrix (batch size $\times$ total features). This operation discards the tensor’s explicit N-D structure. We aim to define a linear transformation that operates directly on N-D tensors, thereby preserving their inherent multi-dimensional organization.
2.2 The NdLinear Transformation
The NdLinear transformation processes N-D tensors by applying separate linear transformations along each of their feature dimensions sequentially. This contrasts with standard linear layers, which flatten the N-D tensor into a 2D matrix, losing the original multi-dimensional structure. By operating on each dimension independently while preserving the others, NdLinear retains the inherent structural information of the data throughout the transformation.
Conceptually, for an input tensor $X\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$, NdLinear learns $N$ distinct weight matrices $W_{1},\ldots,W_{N}$, where each $W_{k}\in\mathbb{R}^{D_{k}\times H_{k}}$ transforms the $k$-th dimension of the input from its original size $D_{k}$ to a new size $H_{k}$, with optional bias vectors $b_{k}\in\mathbb{R}^{H_{k}}$ per dimension. The transformation is applied iteratively: the output from transforming dimension $k$ becomes the input for transforming dimension $k+1$. The procedure is detailed in Algorithm 1.
Algorithm 1 NdLinear Transformation
0:  Input tensor $X\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$ (batch size $B$, original feature dimensions $D_{1},\ldots,D_{N}$), Target output feature dimensions $H_{1},\ldots,H_{N}$, Learnable weight matrices $W_{k}\in\mathbb{R}^{D_{k}\times H_{k}}$ for $k=1,\ldots,N$, Optional learnable bias vectors $b_{k}\in\mathbb{R}^{H_{k}}$ for $k=1,\ldots,N$.
  Let $X_{\text{out}}\leftarrow X$.
  for $k=1$ to $N$ do
     Let current shape of $X_{\text{out}}$ be $(B,S_{1},\ldots,S_{N})$, where $S_{j}=H_{j}$ for $j<k$, and $S_{j}=D_{j}$ else.
     The $k$-th feature dimension (original size $D_{k}$, current size $S_{k}=D_{k}$) is targeted for transformation.
     Permute $X_{\text{out}}$ to move its $k$-th feature dimension to the last position. Shape becomes $(B,S_{1},\ldots,S_{k-1},S_{k+1},\ldots,S_{N},D_{k})$.
     Reshape $X_{\text{out}}$ to a 2D matrix $X_{\text{matrix}}$ of shape $\left(B\cdot\prod_{j\neq k}S_{j},D_{k}\right)$.
     Apply linear transformation: $Y_{\text{matrix}}\leftarrow X_{\text{matrix}}W_{k}+b_{k}$. ($Y_{\text{matrix}}$ has shape $\left(B\cdot\prod_{j\neq k}S_{j},H_{k}\right)$).
     Reshape $Y_{\text{matrix}}$ back to N-D form: $(B,S_{1},\ldots,S_{k-1},S_{k+1},\ldots,S_{N},H_{k})$.
     Permute dimensions to place $H_{k}$ (the new size of the $k$-th feature dimension) back into the $k$-th feature position. Shape becomes $(B,S_{1},\ldots,S_{k-1},H_{k},S_{k+1},\ldots,S_{N})$.
     Update $X_{\text{out}}$ with the result of this transformation.
  end for
  return  $X_{\text{out}}$, now of shape $\mathbb{R}^{B\times H_{1}\times\cdots\times H_{N}}$.
In practice, these operations (transposing, reshaping, matrix multiplication, then inverse reshaping and transposing) can be efficiently implemented using standard tensor library functions like torch.tensordot or einsum. The key is that each weight matrix $W_{k}$ only transforms dimension $D_{k}$ to $H_{k}$. This operation modifies all entries along the $k$-th mode, performing the same linear transformation on each mode-$k$ fiber of the tensor.
This sequential application can be expressed using tensor notation as a series of mode-$k$ tensor-matrix products (Kolda and Bader, 2009): $Y=X\times_{1}W_{1}\times_{2}W_{2}\cdots\times_{N}W_{N}$ , where each product $X\times_{k}W_{k}$ transforms the $k$-th mode of the tensor using matrix $W_{k}$. (Biases $b_{k}$ are added after each mode-$k$ product). The intermediate result of $X\times_{k}W_{k}$ becomes the input for the ×$\times_{k+1}W_{k+1}$ product.
2.3 Relationship to Existing Methods
NdLinear’s design as a structure-preserving, parameter-efficient linear layer for N-D tensors is informed by, yet distinct from, several established concepts in machine learning and tensor algebra.
First, NdLinear’s core mechanism, which is sequential matrix multiplications along each tensor mode, employs fundamental operations from multilinear algebra, prominent in tensor decompositions like Tucker decomposition (Tucker, 1966; Kolda and Bader, 2009). However, unlike classical tensor decomposition methods that primarily analyze or compress static data or pre-existing weight tensors (e.g., (Novikov et al., 2015), (Newman et al., 2024)), NdLinear integrates these mode-wise operations as a learnable, dynamic layer within a network. Its purpose is the efficient forward transformation of activations while preserving N-D structure, not data analysis or post-hoc model compression.
Second, the principle of factorizing operations along data axes for efficiency and structured inductive bias is shared with successful neural network components like depthwise separable convolutions (Chollet, 2017b; Howard et al., 2017b) (separating spatial/channel operations) and axial attention mechanisms (Ho et al., 2019; Wang et al., 2020; Yan et al., 2023) (processing sequence dimensions independently). NdLinear generalizes this factorization: it applies to any N-D tensor and factorizes the linear transformation along all of its modes, not being confined to specific mechanisms or predefined splits (e.g., spatial vs. channel).
Finally, NdLinear belongs to the family of structured linear layers designed for parameter efficiency (Denil et al., 2013), (Wei et al., 2024). While many approaches impose specific algebraic structures on a large conceptual weight matrix (e.g., low-rank (Sainath et al., 2013), or Monarch (Dao et al., 2022) and other forms (Sindhwani et al., 2015), (Potapczynski et al., 2024)), NdLinear’s structure is directly and interpretably derived from the N-D organization of the input data. Factorization occurs along the data’s native modes, offering a data-centric approach to efficient N-D linear transformation.
In essence, NdLinear synthesizes these inspirations into a versatile layer that directly applies linear transformations to N-D data without destructive flattening, prioritizing structural integrity and computational efficiency.
Comparison to Tensor Contraction and Tensor Regression Layers. Tensor Contraction Layers (TCLs) (Cichocki, 2014; Novikov et al., 2015) and Tensor Regression Layers (TRLs) (Kossaifi et al., 2020) also leverage tensor operations. TCLs often focus on dimensionality reduction by contracting specific modes. TRLs extend regression to multi-dimensional data using tensor formats. In contrast, NdLinear aims to transform N-D tensors while fully maintaining their N-D structure throughout the process, rather than primarily for dimensionality reduction or regression. NdLinear leverages the data’s intrinsic N-D nature for parameter efficiency, offering a complementary approach.
3 Analysis of NdLinear: Efficiency and Representational Properties
NdLinear’s design (Section 2) yields significant efficiency advantages and beneficial inductive biases for N-D data. Here, we quantify its parameter and computational efficiency against standard linear layers and examine how its structure shapes representational capacity and guides learning.
3.1 Drastic Parameter Reduction
A hallmark of NdLinear is its exceptional parameter economy. Consider transforming an input tensor $X\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$ to an output $Y^{\prime}\in\mathbb{R}^{B\times H_{1}\times\cdots\times H_{N}}$. A conventional linear layer operating on flattened tensors (input size $D_{\text{flat}}=\prod D_{k}$, output size $H_{\text{flat}}=\prod H_{k}$) requires $D_{\text{flat}}\cdot H_{\text{flat}}+H_{\text{flat}}$ parameters (weights and biases).
NdLinear circumvents this combinatorial growth. By employing $N$ mode-specific transformations, each with a weight matrix $W_{k}\in\mathbb{R}^{D_{k}\times H_{k}}$ and optional bias $b_{k}\in\mathbb{R}^{H_{k}}$, NdLinear’s total parameter count is only $\sum_{k=1}^{N}(D_{k}H_{k}+H_{k})$.
The practical impact is profound. For example, transforming a 3D tensor where all $D_{k}=H_{k}=D=32$, a standard linear layer requires $O(D^{2N})$ parameters, which is $32^{6}\approx 10^{9}$ parameters. In contrast, NdLinear requires $O(ND^{2})$ parameters, merely $3\cdot 32^{2}\approx 3\times 10^{3}$, a reduction of several orders of magnitude. This drastic reduction makes NdLinear highly advantageous for leaner models, reduced memory usage, and potentially mitigating overfitting, especially with high-dimensional N-D inputs.
3.2 Efficient Computation
Corresponding to its parameter economy, NdLinear demonstrates significant computational efficiency. A standard linear layer performs approximately $2B\cdot(\prod D_{k})\cdot(\prod H_{k})$ Floating Point Operations (FLOPs).
NdLinear involves $N$ sequential mode-wise transformations. The total FLOPs are:
$\text{FLOPs}_{\text{NdLinear}}\approx 2B\sum_{k=1}^{N}\left[\left(\prod_{j=1}^{k-1}H_{j}\right)\left(\prod_{j=k+1}^{N}D_{j}\right)D_{k}H_{k}\right].$ (1)
This sum is typically far smaller than the product for a standard layer. Using the previous example ($N=3,D_{k}=H_{k}=D=32$), a standard layer requires $O(BD^{2N})$ FLOPs ($\approx B\cdot 10^{9}$). NdLinear requires $O(NBD^{N+1})$ FLOPs ($\approx B\cdot 3\cdot 10^{6}$). This represents orders of magnitude in computational savings, making NdLinear significantly faster, especially as $N$ or dimension sizes $D_{k},H_{k}$ increase. Its sequential operations also allow for efficient memory management of intermediate results.
3.3 Structured Representation and Beneficial Inductive Bias
NdLinear’s efficiency stems directly from its structured, factorized transformation of N-D tensors, a design that inherently shapes its representational capabilities and introduces inductive biases.
A single NdLinear layer, with its mode-wise factorization and reduced parameter set, represents a constrained subset of all possible linear transformations. This constraint is its primary inductive bias: NdLinear is predisposed to learn "separable patterns" aligned with the data’s intrinsic axes through sequential, dimension-wise transformations. This bias is particularly potent for data with inherent modal structures (e.g., temporal, spatial, channel-wise), guiding the model to leverage this organization for more efficient learning and better generalization than possible from flattened vectors.
This structural constraint also induces strong regularization. By limiting model complexity compared to dense layers, NdLinear naturally mitigates overfitting, especially with limited training data. This is akin to benefits from explicit complexity reduction (e.g., low-rank assumptions (Sainath et al., 2013; Denil et al., 2013)), achieved here through NdLinear’s intrinsic low-complexity design.
4 Experimental Results
To empirically validate NdLinear, we conducted comprehensive experiments assessing its parameter efficiency, performance across diverse applications, and scalability. This section details the experimental setups, key results, and discusses the derived insights, organized by application domain.
4.1 Experimental Overview
Our extensive empirical investigation assesses NdLinear’s impact on parameter efficiency, its operational versatility, and its scalability. We benchmark NdLinear by integrating it as a drop-in replacement for standard Linear layers within influential neural architectures across key domains, demonstrating its benefits through a rigorous and wide-ranging suite of tests:
  • Natural Language Processing (Section 4.2): Our NLP evaluation is extensive, covering large language model pretraining (adapting OPT variants (Zhang et al., 2022)), text classification with BERT-based models (Devlin et al., 2019), and parameter-efficient finetuning. For the latter, we introduce and assess NdLinear-LoRA, showcasing its effectiveness on diverse mathematical reasoning and commonsense QA benchmarks with prominent models such as Qwen3-1.7B and LLaMA3-8B.
  • Time Series Analysis (Section 4.3): We investigate NdLinear’s performance in RNN-based models (Hochreiter and Schmidhuber, 1997) and Transformer-based models (Zerveas et al., 2021) for both forecasting and classification tasks. These evaluations span ten distinct datasets from the widely-used ETT and UCR archives.
  • Tabular Data (Section 4.4): We apply NdLinear in MLPs for tabular classification and regression tasks (Shwartz-Ziv and Armon, 2022), demonstrating its utility on this common data modality.
  • Computer Vision (Section 4.5): We evaluate NdLinear across a spectrum of vision tasks, including Vision Transformers (ViTs) (Dosovitskiy et al., 2021) for model distillation, CNNs (Krizhevsky et al., 2012b) for image classification (with detailed weight behavior analysis), and Diffusion Transformers (DiTs) (Peebles and Xie, 2023) for generative modeling.
This diverse set of experiments, encompassing varied architectures and numerous datasets per domain, allows for a thorough assessment of NdLinear. In all evaluations, NdLinear directly substitutes conventional Linear layers in critical network components (e.g., feature extractors, MLP blocks, timestep encoders, projection layers, and low-rank adapters). See more details and our fully reproducible code in the Appendix C.
4.2 Natural Language Processing Applications
4.2.1 Parameter-Efficient Finetuning with LoRA and NdLinear-LoRA
Low-rank adaptation (LoRA) (Hu et al., 2022) is a popular parameter-efficient strategy for finetuning LLMs. We investigate if a NdLinear layer, termed NdLinear-LoRA, can effectively replace the standard LoRA architecture (composed of two rank-$r$ matrices) while maintaining or improving downstream task performance with even fewer trainable parameters.
Setup. We fine-tune two LLMs, Qwen3-1.7B (Yang et al., 2025) and LLaMA3-8B (Dubey et al., 2024). Training utilizes the OpenMathInstruct-1 (math10k) corpus (Toshniwal et al., 2024) for mathematical reasoning and the CommonsenseQA training split (Talmor et al., 2019) for commonsense reasoning.
Results. Table 1 presents accuracy on Math benchmarks (GSM8K (Cobbe et al., 2021), MultiArith (Roy and Roth, 2016)) and Commonsense Reasoning benchmarks (CSQA (Talmor et al., 2019), ARC-easy, ARC-challenge (Clark et al., 2018), BoolQ (Clark et al., 2019)). NdLinear-LoRA generally achieves comparable or superior accuracy to LoRA with significantly fewer trainable parameters (4x to 9x reduction compared to LoRA r=8).
  • Qwen3-1.7B: With only 1.15M trainable parameters (vs. 8.72M for LoRA r=8), NdLinear-LoRA shows substantial gains on GSM8K (+11.9 points over LoRA r=8) and improves performance on MultiArith. It achieves results comparable to LoRA r=8 on CSQA and ARC-e, and slightly trails on ARC-c.
  • LLaMA3-8B: Using only 2.26M parameters (vs. 20.97M for LoRA r=8, a 9$\approx$ 9$\times$ 9 × reduction), NdLinear-LoRA matches or exceeds LoRA r=8 on CSQA and ARC-c, and LoRA r=4 on ARC-e. While performance on GSM8K and MultiArith is lower than LoRA, the results are achieved with considerably fewer parameters.
Discussion.
These experiments, utilizing a straightforward implementation of NdLinear-LoRA, suggest that NdLinear’s structured transformations offer an effective alternative to LoRA’s low-rank matrix factorization. This presents a promising avenue for developing even more memory-efficient adaptation techniques for LLMs without performance degradation.
Table 1: NdLinear-LoRA demonstrates significant parameter efficiency for LLM finetuning, achieving comparable or improved accuracy over standard LoRA methods with up to 9×9$\times$9 × fewer trainable parameters. Accuracy ($\uparrow$) after LoRA of Qwen3-1.7B and LLaMA3-8B models.
Math CS Reasoning
Model Method Params GSM8K M.Arith CSQA ARC-e ARC-c BoolQ
Qwen3-1.7B LoRA ($r{=}4$) 4.36M 45.6 88.9 80.4 91.9 79.4 79.7
LoRA ($r{=}8$) 8.72M 40.3 82.2 80.9 91.8 79.3 80.8
NdLinear-LoRA 1.15M 52.2 90.0 81.0 92.2 78.3 79.7
LLaMA3-8B LoRA ($r{=}4$) 10.48M 50.5 84.4 80.6 90.4 76.3 85.1
LoRA ($r{=}8$) 20.97M 51.6 81.1 81.7 89.0 73.6 76.5
NdLinear-LoRA 2.26M 40.2 80.0 82.9 90.9 76.6 80.5
4.2.2 Efficient Large Language Model Pretraining
To evaluate NdLinear’s effectiveness at scale, we adapted the Open Pretrained Transformer (OPT) architecture (Zhang et al., 2022) by replacing Linear layers with NdLinear in the feedforward blocks.
Setup. We evaluated two variants: OPT-Small (124M parameters) and OPT-Mid (350M parameters). For these proof-of-concept experiments, both were trained from scratch using a subset of the original training corpus (BookCorpus and Wiki40B-English).
Results. Despite having fewer parameters, both OPT models with NdLinear layers achieved lower perplexity (Table 4 in Appendix D). The improvement was more pronounced in the larger model (perplexity gap increased from 0.215 in OPT-Small to 0.361 in OPT-Mid), suggesting that NdLinear’s benefits may scale with model size. The training and evaluation loss curves (Figure 1) further validate these improvements, showing consistently lower loss for NdLinear variants.
4.2.3 Efficient Text Classification with BERT
We evaluated NdLinear for text classification using BERT (Devlin et al., 2019) on the SST-2 (Socher et al., 2013) and CoLA (Warstadt et al., 2018) datasets.
Setup. In the BERT model, we replaced the conventional two-layer Linear classification head with a single NdLinear layer followed by a standard Linear classification layer.
Results. As shown in Table 6 in Appendix D , NdLinear improved both accuracy and ROC AUC on both datasets while reducing the parameter count in the classification head by approximately 85%.
4.3 Time Series Analysis Applications
We evaluated NdLinear’s effectiveness in time series analysis, focusing on multivariate forecasting with RNNs and Transformers, and on classification tasks with Transformer-based models.
4.3.1 Efficient Multivariate Time Series Forecasting
We tested NdLinear in RNNs and Transformers for forecasting on four ETT (Electricity Transformer Temperature) datasets (Zhou et al., 2021).
Setup. The task was to predict the next 12 hours of data based on 24 hours of historical input. For RNNs, we compared a standard RNN against an RNN with NdLinear. For Transformers, we compared a 1-layer Transformer against a Transformer where Linear layers in the FeedForward blocks were replaced by NdLinear layers.
Results. NdLinear consistently improved forecasting accuracy (lower MSE and MAE) across all datasets for both RNNs and Transformers (Table 3), while using significantly fewer parameters. The improvements were particularly notable on the ETTh1 and ETTh2 datasets.
4.3.2 Efficient Time Series Classification
Transformer models, while successful for time series classification, often have high computational demands due to their large parameter counts. We investigated NdLinear’s potential to mitigate this by enhancing parameter efficiency while maintaining or improving performance.
Setup. We applied Transformer-based models to six time series classification datasets from the UCR Archive (Dau et al., 2019). We replaced the standard Linear layers in the Transformer blocks with NdLinear layers. For a fair comparison, we also included a smaller standard Linear variant (denoted Base Model) where its hidden dimension matches the first transformation dimension of the corresponding NdLinear layer. F1 scores are reported in Table 7 in Appendix D.
Table 2: NdLinear significantly enhances Transformer-based time series classification, achieving superior F1 scores across UCR datasets with comparable or substantially fewer parameters than baseline models. F1 Score and efficiency (# Params) of NdLinear compared to a standard Base Model and a parameter-comparable Base Model (whose parameters are aligned with NdLinear’s).
Model Metric ECGFiveDay HeartBeat Chlorine Conc. ECG5000 LSST Sleep
Base Model Params 3 363 4 323 12 900 12 966 29 343 12 966
F1 Score 0.7668 0.7250 0.4436 0.8886 0.5928 0.4911
Base Model Params 1 779 2 739 6 660 6 726 15 375 6 726
F1 Score 0.7624 0.7214 0.4440 0.8878 0.5907 0.4897
NdLinear Params 1 804 2 752 6 709 6 823 15 472 6 823
F1 Score 0.7783 0.7363 0.4571 0.9058 0.6486 0.4978
Table 3: NdLinear substantially improves time-series forecasting accuracy and dramatically reduces parameter counts in both RNN and Transformer models on ETT datasets. The performance comparison on four ETT (Electricity Transformer Temperature) datasets. With NdLinear, model parameters are reduced by 50%, alongside consistent gains in accuracy. Metrics reported are mean $\pm$ standard deviation over three experimental runs.
RNN Transformer
Dataset Method (Params: RNN / Trans.) MSE MAE MSE MAE
ETTh1 Linear (20.5k / 138k) 0.2900 $\pm$ 0.0170 0.4060 $\pm$ 0.0246 0.0217 $\pm$ 0.0001 0.1158 $\pm$ 0.0004
NdLinear (9.6k / 70k) 0.0880 $\pm$ 0.0115 0.2204 $\pm$ 0.0105 0.0173 $\pm$ 0.0003 0.0995 $\pm$ 0.0005
ETTh2 Linear (20.5k / 138k) 0.2636 $\pm$ 0.0949 0.3955 $\pm$ 0.0748 0.0226 $\pm$ 0.0001 0.1229 $\pm$ 0.0003
NdLinear (9.6k / 70k) 0.1536 $\pm$ 0.0137 0.2831 $\pm$ 0.0119 0.0158 $\pm$ 0.0019 0.0995 $\pm$ 0.0071
ETTm1 Linear (20.5k / 138k) 0.0187 $\pm$ 0.0012 0.0926 $\pm$ 0.0039 0.0180 $\pm$ 0.0001 0.1005 $\pm$ 0.0005
NdLinear (9.6k / 70k) 0.0174 $\pm$ 0.0017 0.0894 $\pm$ 0.0039 0.0161 $\pm$ 0.0006 0.0936 $\pm$ 0.0027
ETTm2 Linear (20.5k / 138k) 0.0148 $\pm$ 0.0007 0.0825 $\pm$ 0.0047 0.0151 $\pm$ 0.0000 0.0965 $\pm$ 0.0001
NdLinear (9.6k / 70k) 0.0139 $\pm$ 0.0009 0.0797 $\pm$ 0.0039 0.0141 $\pm$ 0.0001 0.0929 $\pm$ 0.0004
Results. NdLinear reduces parameter counts by up to 47% compared to the Linear layers in Transformers, while consistently improving F1 scores across all six datasets. Notably, even the standard Linear model with nearly twice the parameters often lags behind the NdLinear variant, highlighting NdLinear’s superior parameter efficiency and representational capacity for these tasks. The Linear variant, despite having a similar number of parameters to NdLinear, generally underperforms NdLinear, further emphasizing the architectural benefits of the NdLinear approach.
4.4 Tabular Data Applications
We evaluated NdLinear’s performance with MLPs on tabular data for both classification and regression tasks, demonstrating its ability to improve efficiency and performance in this common data modality.
Setup. We designed an MLP comprising two feature-extraction layers followed by a final linear layer for prediction. Within the feature-extraction layers, we compared blocks using standard Linear layers against blocks using NdLinear layers. Experiments were conducted on the Cardio Disease dataset (Sulianova, 2025) for classification and the Food Delivery Time dataset (Kumar, 2025) for regression.
Results. As shown in Table 7 in Appendix D, NdLinear demonstrated superior performance on both datasets while significantly reducing the number of parameters. For the Cardio Disease classification task, NdLinear achieved higher accuracy (0.7321 vs. 0.7265) with 67% fewer parameters. In the Food Delivery Time regression task, NdLinear yielded a lower MSE (67.824 vs. 70.508) with 58% fewer parameters.
4.5 Computer Vision Applications
We evaluated NdLinear in several computer vision tasks, including image classification with both ViTs and CNNs, and generative tasks with DiTs, to demonstrate its parameter efficiency and performance.
4.5.1 Parameter-Efficient Vision Transformer Distillation
Setup. We explore NdLinear in ViTs (Dosovitskiy et al., 2021) for model distillation. A pre-trained ViT-B/16 model from (PyTorch, ) serves as the teacher. Student models, termed NdViT, are created by replacing all Linear layers in the feedforward blocks of a standard ViT with NdLinear layers. We experiment with various configurations, altering student model dimensions and the number of transformer blocks, evaluating on CIFAR-10 and CIFAR-100.
Results. NdViT consistently outperforms conventional ViT baselines while using 26-68% fewer parameters (Table 9 and Figure 2 in Appendix D). Notably, an NdViT with embedding dimension 200 surpasses a naive ViT with dimension 500, highlighting NdLinear’s parameter efficiency. The performance advantage of NdViT tends to increase with more transformer blocks and larger dimensions.
4.5.2 Efficient Image Classification with CNNs
Setup. We trained a CNN on CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009). The architecture consists of two convolutional layers followed by either a standard Linear layer or an NdLinear layer, and one final Linear layer as the classification head.
Results. NdLinear achieves higher accuracy with significantly fewer parameters on both datasets (Table 8 in Appendix D). On CIFAR-10, NdLinear improves accuracy by 2.6% while reducing parameters by 94%. On CIFAR-100, it improves accuracy by 5.1% with 60% fewer parameters.
4.5.3 Efficient Generative Modeling with Diffusion Transformers (DiT)
We evaluated NdLinear in DiTs by training models from scratch on ImageNet-100 (Russakovsky et al., 2015) and by modifying pre-trained DiT models from (Jin and Xie, 2024). For image generation, FID (Dhariwal and Nichol, 2021) was the evaluation metric.
Training DiT from Scratch
To assess NdLinear’s effectiveness when training from scratch, we aimed to achieve superior performance with a comparable number of parameters. We replaced the two Linear layers in the timestep embedder of the standard DiT with NdLinear layers.
Results. NdLinear variants consistently achieve lower (better) FID-10k scores compared to baseline DiTs when trained from scratch with similar parameter counts (Figure 3 in Appendix D).
Modifying Linear Layers in Pre-trained DiT Models
To evaluate performance maintenance with parameter reduction, we replaced standard Linear layers with NdLinear in pre-trained DiTs. This was done in both the timestep embedder and the MLP components of attention heads.
Results. Table 10 in Appendix D shows that NdLinear-based models (619M and 563M) achieve FID scores comparable to the larger baseline 674M model, despite significant parameter reductions.
5 Conclusion
This paper introduced NdLinear, a novel linear transformation layer meticulously designed to address the prevalent issue of destructive input flattening. By operating directly on N-D tensors and applying transformations sequentially along each dimension, NdLinear fundamentally preserves the rich, inherent structural integrity of multi-dimensional data. This offers a robust alternative to conventional linear layers that often discard critical cross-axis information, thereby enabling a more faithful representation of N-D data throughout neural network processing.
Our comprehensive experiments across diverse domains as computer vision, natural language processing, time series, and tabular data, consistently demonstrated NdLinear’s compelling advantages. It achieves substantial parameter efficiency, with reductions ranging from valuable percentages in transformer blocks to over 90% in contexts like CNN feature extractors or compact classification heads. This efficiency is achieved while typically maintaining or, crucially, enhancing model performance. These improvements, observed across a variety of architectures (CNNs, RNNs, Transformers, MLPs), underscore NdLinear’s generality as a drop-in replacement. These empirical results strongly affirm our analysis regarding NdLinear’s beneficial inductive bias and efficiency gains.
By embedding N-D structure awareness into a foundational layer, NdLinear encourages a paradigm shift in designing neural architectures, moving from attempts to recover lost data relationships towards actively preserving them. This principled approach makes NdLinear a valuable component for building more capable and efficient large-scale AI systems, potentially reducing the need for certain compensatory preprocessing steps. The demonstrated strengths in efficiency and structure-aware representation learning open promising avenues in computationally intensive domains such as generative AI, complex scientific discovery (e.g., in genomics and multimodal data analysis), and in developing richer representations for intelligent agents, all of which warrant further exploration.
In summary, NdLinear provides a simple yet potent modification to a ubiquitous component of neural networks. By championing the preservation of N-D data structures as its core principle—to process data in its native form rather than flatten it—NdLinear offers a foundational advance towards more efficient, structurally-aware, and ultimately more effective AI models capable of better navigating our multi-dimensional world.
Appendix A Comparison with Structured Representation Learning Methods
We now compare NdLinear to other approaches that aim to capture structured dependencies or reduce parameters in high-dimensional models:
Graph Neural Networks (GNNs).
GNNs handle graph-structured data [Scarselli et al., 2008, Micheli, 2009, Bronstein et al., 2017, Zhou et al., 2020, Wu et al., 2020]. One can represent an $n$-dimensional grid (e.g., an image) as a graph, with each element as a node connected to its neighbors. GNNs often use graph convolutions or message passing [Gilmer et al., 2017, Kipf and Welling, 2017, Hamilton et al., 2017], capturing local structure without flattening. They aggregate features from neighboring nodes, acting like a convolution or local smoothing.
However, using GNNs for grid data can be overkill. They add significant computational overhead and do not leverage the regular structure. To reach global interactions, GNNs usually need many layers or long-range edges [Battaglia et al., 2018]. By contrast, NdLinear applies a global linear interaction along each dimension in one layer. A GNN typically shares weights across edges or edge types, so parameters do not necessarily grow with input size. Still, it may require many layers to propagate information globally. NdLinear instead uses the grid’s structure to apply dimension-wise interactions in one step, making it simpler and more efficient for tensor grids. GNNs excel at arbitrary graphs where no regular structure exists. Another difference is weight sharing: NdLinear shares weights fully along each dimension, while GNNs can have more complex or no sharing.
In summary, NdLinear is specialized for tensor grids and is faster and simpler on those tasks, whereas GNNs are more flexible but heavier. Hence, one typically does not replace a fully-connected layer for images with a GNN, due to the overhead.
Tensor Decomposition Methods.
NdLinear is closely related to tensor decomposition methods used for model compression. Examples include Tucker [Tucker, 1966], CP (CANDECOMP/PARAFAC) [Carroll and Chang, 1970, Harshman, 1970], and Tensor-Train (TT) Decompositions [Oseledets, 2011, Novikov et al., 2015]. These methods view a large weight matrix (like in fully-connected layers) as a high-order tensor. They factorize this tensor into smaller components. For instance, the Tensor-Train decomposition represents weights as sequences of smaller tensors. This achieves significant parameter reduction with minimal accuracy loss [Novikov et al., 2015].
NdLinear can be seen as a hand-crafted factorization of a fully-connected weight matrix. Specifically, the full weight matrix $W_{\text{full}}$ implicitly has a Kronecker product structure derived from mode-wise matrices $\{W_{1},\dots,W_{n}\}$. This corresponds to a rank-1 Tucker decomposition without a core tensor (or equivalently, a core of rank 1 in each mode).
The main trade-off here is expressiveness vs. efficiency. NdLinear’s decomposition is low-rank in a multilinear sense. It cannot represent unique, non-factorizable interactions between dimensions. More flexible decompositions, such as full Tucker or higher-rank tensor decompositions, capture more interactions. However, these approaches require significantly more parameters compared to NdLinear’s simple sum $\sum_{i}D_{i}H_{i}$. They can also be more challenging to train, sometimes needing special initialization or multi-stage training.
To combat these, NdLinear balances these trade-offs well. It is highly parameter-efficient, easy to train from scratch, and captures essential structured dependencies. If needed, one can extend NdLinear by increasing the factorization rank. For example, one could learn multiple $W_{i}$ matrices per mode and sum their effects, analogous to a rank-$R$ core. However, our experiments show that the simple version already performs very well. Compared to general tensor decomposition methods, NdLinear is also more interpretable. Each $W_{i}$ clearly indicates how dimension $i$ is transformed, rather than dispersing transformations across multiple factors.
Grouped and Factorized Convolutions.
Various convolutional factorizations reduce CNN computation, notably:
  • Grouped convolutions [Krizhevsky et al., 2012a, Xie et al., 2017] restrict filters to subsets of channels, reducing parameters.
  • Depthwise separable convolutions [Howard et al., 2017a, Chollet, 2017a] split convolution into per-channel depthwise and pointwise ($1\times 1$) operations, significantly cutting FLOPs.
NdLinear applies a similar concept to fully-connected layers, decomposing weights by dimension rather than channel. Instead of one large weight matrix mapping $\prod D_{i}$ inputs to $\prod H_{i}$ outputs, NdLinear factorizes it into $n$ matrices, each mapping $D_{i}$ to $H_{i}$. This sequential dimension-wise grouping yields structured weights, improving efficiency with minimal accuracy loss.
Unlike grouped or separable convolutions that typically serve as internal layers, NdLinear naturally produces structured multi-dimensional outputs $(H_{1},\dots,H_{n})$, making it suitable for structured prediction tasks.
Other Specialized Layers.
Several specialized approaches for structured representation learning have been explored:
  • Slicing-based layers [Shao et al., 2016, Dieleman et al., 2016] separately process spatial, temporal, or rotated input segments, preserving structure without flattening.
  • Capsule Networks [Hinton et al., 2011, Sabour et al., 2017, Hinton et al., 2018] use vector or matrix capsules with routing mechanisms to maintain hierarchical relationships.
  • Hadamard/Fourier feature mixing [Rahimi and Recht, 2007, Le et al., 2013, Tancik et al., 2020, Lee-Thorp et al., 2022, Pan et al., 2022] employs fixed, non-learnable transforms (e.g., Walsh-Hadamard, Fourier) for global mixing.
NdLinear is simpler and more flexible than these methods, applying learned factorized linear transformations along each dimension without specialized routing or static transforms.
Appendix B More Related Work
Deep neural networks often contain millions of parameters, many of which are highly redundant. For instance, Denil et al. [2013] demonstrate that a large fraction of weights can be predicted from a small subset. In some cases, 95% of parameters need not be learned with no loss in accuracy. This redundancy has motivated numerous efficient parameterization strategies that preserve model expressiveness while drastically reducing storage and computation requirements.
Structured Tensor Factorization.
A prominent research direction exploits high-order (multi-dimensional) structures of weight tensors using tensor factorization. Lebedev et al. [2015] apply CP decomposition to 4D convolutional kernels, reducing parameters and inference cost. Novikov et al. [2015] introduce Tensor Train (TT) layers, compressing fully-connected layers into compact tensors without losing expressiveness. More recently, Ye et al. [2020] propose Block-Term Tensor Networks, approximating network weights through block-term (Tucker) decompositions. These tensor-structured layers significantly reduce parameters and preserve high-order interactions.
Structured Matrices and Parameter Sharing.
Other methods impose algebraic structure on weight matrices or use parameter sharing to reduce model size. Sindhwani et al. [2015] replace dense matrices with Toeplitz-like structures, significantly compressing layers without losing performance. Similar methods include circulant matrices, block-circulant transforms, and low-rank factorizations [Lebedev et al., 2015]. These structured layers consistently match or exceed unstructured models in performance, using fewer parameters and lower computational cost.
Multi-Space Representations.
Emerging multi-space learning methods embed data into multiple latent spaces to preserve complex relationships. Rather than relying only on Euclidean spaces, these methods project data into multiple spaces (e.g., Euclidean and hyperbolic) to capture richer structures. Wang et al. [2023] embed LiDAR features in Euclidean and hyperbolic spaces, improving hierarchical information encoding and pose regression. Multi-space learning complements tensor factorization by enhancing model expressiveness without increasing layer size.
Preserving High-Order Structure in Practice.
A common theme across these efforts is that preserving high-order structure leads to compact yet expressive representations. Viewing weight or activation tensors in their natural multi-dimensional format directly captures interactions along each dimension. Depthwise separable convolutions [Chollet, 2017b, Howard et al., 2017b] partially achieve this by separating channel-wise and spatial-wise operations. However, most fully-connected layers still flatten inputs, ignoring multi-axis dependencies.
The proposed NdLinear approach naturally fits within structured methods. NdLinear iterates through each dimension of an $n$-dimensional tensor, applying dimension-specific linear transforms. It preserves multi-axis relationships without extra computational cost. Similar to depthwise separable convolutions, NdLinear generalizes efficiently to more dimensions and tasks beyond convolutional features. Combining NdLinear with existing factorization methods can further enhance multi-space learning in large-scale models.
These advances in structured tensor methods, multi-space feature learning, and efficient parameterization show that deep networks can be both lightweight and expressive. Exploiting algebraic structures or multiple embedding spaces yields compact models, often with improved accuracy—essential for large-scale or resource-constrained applications.
Appendix C Implementation Details and Training Protocol
Here we present implementation details and training protocol of NdLinear (Algorithm 1). Implementing NdLinear in practice involves careful attention to efficiency and compatibility with existing deep learning frameworks. We outline key considerations in the following.
Memory Efficiency. Despite handling high-dimensional tensors, NdLinear is memory-efficient due to its factorized parameterization. The forward pass requires allocating intermediate tensors during each mode transformation (after each linear operation, the tensor has one updated dimension). However, these intermediate allocations are of the same order as the input/output size and significantly smaller than the memory required for a gigantic flattened weight matrix. Modern tensor libraries (PyTorch, TensorFlow) facilitate implementing transpose-reshape-multiply steps without excessive data copying. We ensure in-place operations where possible (e.g., using $\mathtt{view}$ in PyTorch). Operations primarily reuse the input buffer for output as each mode is transformed, ensuring modest peak memory usage.
Parameter Initialization. Each weight matrix $W_{i}$ can be initialized using standard strategies for linear layers (Xavier/Glorot [Glorot and Bengio, 2010] or Kaiming [He et al., 2015] initialization based on fan-in and fan-out). Since $W_{i}$ has fan-in $D_{i}$ and fan-out $H_{i}$, the initialization follows ${\rm Uni}(-\sqrt{\frac{6}{D_{i}+H_{i}}},\sqrt{\frac{6}{D_{i}+H_{i}}})$ for Xavier uniform, or analogous formulas for other initializations. This helps maintain stable gradients across modes. One subtle point is that if $n$ is large, each mode’s weight is relatively small, mitigating the risk of extremely large fan-in. We observed no initialization-specific difficulties; indeed, NdLinear’s parameter reduction may help avoid gradient explosion or vanishing issues in deep networks.
Computational Overhead. Factorized operations (multiple transpose and reshape operations with smaller matrix multiplications) are highly optimized in modern BLAS libraries. Practically, runtime is comparable to or faster than fully-connected layers with similar outputs, due to reduced total FLOPs. Python-level overhead is minimal; the algorithm can be implemented in a single forward function looping over modes. For moderate $n$ (up to 4 or 5 dimensions), this loop is short. Explicit loops or unrolling (NdLinear2d, NdLinear3d, etc.) are feasible, but a simple loop suffices. Autograd engines handle tensor operations seamlessly, allowing standard backpropagation. Each $W_{i}$ receives gradients normally from upstream gradients.
Training Protocols. NdLinear layers can be trained end-to-end with standard optimization algorithms (SGD, Adam) just like standard linear layers. Loss functions depend on the task (cross-entropy, MSE, etc.) and are unaffected by NdLinear. However, because NdLinear significantly reduces parameters compared to fully-connected layers, it tends to overfit less, possibly needing less aggressive regularization. Common techniques remain useful: weight decay (L2 regularization) on weights $W_{i}$, and optionally dropout between layers. Dropout can be applied before or after NdLinear; entries in the output tensor $Y$ can be dropped as usual. Specialized regularizers for factorized weights (norm regularization, orthogonality) may help further restrict solution spaces, though not required.
Optimization and Convergence. Practically, each $W_{i}$ is updated based on a portion of the overall error gradient (due to sequential mode transforms). In experiments, all $W_{i}$ matrices learned smoothly with default optimizer settings. If dimensionality varies significantly across modes, gradient clipping or adaptive learning rates per mode may be beneficial. Throughout our numerical investigations, training dynamics are stable overall — NdLinear layers integrated seamlessly into models without requiring special tuning. Standard protocols (learning rate schedules, early stopping criteria, etc.) used for equivalent models with dense layers apply here.
Appendix D Full Experimental Details and Results
We present full experimental details and more results on LoRA fine-tuning in section D.1; language-model pretraining (OPT and BERT) in section D.1; time-series prediction (RNN and Transformer) in section D.2; tabular data in section D.3; and vision tasks (CNN, ViT, and DiT) in section D.4. The complete experiment code is available at https://github.com/cyclone-trout/ndlinear_neurips.
D.1 Parameter-Efficient Finetuning with LoRA and NdLinear-LoRA
In our study, we utilized state-of-the-art transformer architectures to investigate the impact of targeted modu le adaptation. We selected two base models, Qwen3-1.7B-Base [Yang et al., 2025] and Meta-Llama-3-8B [Dubey et al., 2024], recognized for their robust performance across various tasks. To focus our adaptations, we targeted specific modules within these models, including q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj. This approach allowed us to enhance model capacity and efficiency selectively.
For the adaptation process, we employed Low-Rank Adaptation (LoRA) techniques [Hu et al., 2022], specifically using both NdLinear LoRA and classic LoRA configurations. We explored a range of alpha values (1, 4, and 8) and rank settings (1, 4, and 8) to determine the optimal configuration for our models. This exploration was critical for understanding how different levels of parameter sharing and scaling affect model performance and generalization.
The training process was conducted using the AdamW optimizer, a choice informed by its effectiveness in managing the complexities of transformer models. We set the learning rate to $1\times 10^{-4}$, which provided a suitable balance between convergence speed and training stability. The batch size was set to 1, a decision that facilitated the use of gradient accumulation to optimize GPU memory usage. To ensure the models could handle a wide variety of inputs, we set the maximum sequence length to 512 tokens. The models were trained over 2 epochs, a duration found to be sufficient for achieving significant performance improvements without excessive computational cost. To ensure reproducibility, we used a random seed of 42 across all experiments.
Our models were fine-tuned using two datasets: Math10K and CommonsenseQA. These datasets were chosen for their ability to challenge the models with both mathematical reasoning and commonsense understanding. For evaluation, we employed a diverse set of benchmark datasets, including GSM8K, MultiArith, ARC-C, ARC-E, and BoolQ. This selection allowed us to assess the models’ generalization capabilities across different types of reasoning tasks.
The entire implementation was carried out on a single NVIDIA H100 GPU, using Hugging Face’s AutoModelForCausalLM framework [^1] https://huggingface.co/docs/transformers/en/model_doc/auto, integrated with our custom NdLinear adapter layer. Datasets were tokenized using the default tokenizer for each model, with padding applied to the eos_token. We employed label masking to exclude prompt tokens from loss computation, ensuring that training focused on the relevant portions of the input. Our implementation leveraged PyTorch, along with Hugging Face Transformers, PEFT, and Accelerate, to facilitate efficient model training and adaptation. Evaluation was performed in a zero-shot setting using greedy decoding, which provided a consistent measure of model performance without the variability introduced by sampling methods.
Open Pretrained Transformer (OPT) [Zhang et al., 2022]. For OPT-Small, which originally contains 124M parameters, replacing the standard linear layers with NdLinear reduces the parameter count to 119M. Similarly, for OPT-Mid, the parameter count decreases from 350M to 337M after the substitution.
Table 4: Perplexity comparison for OPT-Small and OPT-Mid models with Linear vs. NdLinear layers.
Linear NdLinear
OPT-Small (Params) 15.970 (124M) 15.755 (119M)
OPT-Mid (Params) 12.926 (350M) 12.565 (337M)
In Table 4, both the OPT-Small and OPT-Mid models achieve lower perplexity scores after replacing standard linear layers with NdLinear layers, despite having fewer parameters. Moreover, the performance improvement becomes more significant as model size increases, with the perplexity gap widening from $0.215$ in OPT-Small to $0.361$ in OPT-Mid. Figure 1 shows that OPT models with NdLinear feedforward layers achieve lower final training and evaluation losses compared to their counterparts using standard linear feedforward layers.
(a) Training Loss during Pre-Training.
(b) Evaluation Loss during Pre-Training.
Figure 1: Training and evaluation loss curves during OPT model pretraining. NdLinear variants consistently achieve lower loss values.
Zero-shot Tasks. We also evaluate the OPTs’ pretraining on 10 zero-shot NLP tasks:
  • Natural Language Inference Tasks: CB [De Marneffe et al., 2019]
  • Coreference Resolution Tasks: Winogrande [Sakaguchi et al., 2021]
  • Sentence Completion Tasks: COPA [Roemmele et al., 2011], HellaSwag [Zellers et al., 2019]
  • Word Sense Disambiguation Tasks: WiC [Pilehvar and Camacho-Collados, 2018]
  • Question Answering Tasks: ARC-Easy, ARC-Challenge [Clark et al., 2018], OpenBookQA [Mihaylov et al., 2018], BoolQ [Clark et al., 2019]
  • Commonsense Reasoning Tasks: PIQA [Bisk et al., 2020]
During evaluation, we cast all of the above tasks into a multiple-choice format. Namely, the goal is to select the correct completion from a set of candidate options. For each option, we compute the language model (LM) likelihood of the full input consisting of the context concatenated with the candidate completion. To account for differences in the lengths of candidate completions, we compute the average per-token log-likelihood for each option, following [Brown et al., 2020]. The model’s prediction is taken to be the option with the highest per-token likelihood.
Table 5: Perplexity Score and Zero-Shot Performance on OPT Model with and without NdLinear.
OPT-Small OPT-Mid
     Linear NdLinear      Linear NdLinear
Num of Params 124M 119M 350M 337M
Perplexity 15.970 15.755 12.926 12.565
CB 0.32 0.38 0.50 0.52
Winogrande 0.49 0.51 0.50 0.50
COPA 0.58 0.53 0.54 0.56
HellaSwag 0.26 0.26 0.28 0.29
WiC 0.51 0.50 0.50 0.49
ARC-Easy 0.29 0.30 0.29 0.29
ARC-Challenge 0.24 0.25 0.24 0.23
OpenBookQA 0.32 0.35 0.34 0.34
BoolQ 0.49 0.59 0.60 0.56
PIQA 0.53 0.50 0.53 0.53
BERT [Devlin et al., 2019]. We replace the conventional two-layer linear classification head in BERT with an NdLinear layer followed by a classification layer. The NdLinear transforms have hidden dimensions of $(2,2)$. Each model is trained for 200 epochs with a batch size of 32, a hidden layer size of 128, and a learning rate of 0.005.
Table 6: BERT text classification performance on CoLA and SST-2 datasets. NdLinear improves accuracy and ROC AUC with $\approx$85fewer parameters in the classification head.
Dataset Method Params (Head) Accuracy ROC AUC
CoLA Linear 1,544 0.7790 $\pm$ 0.0143 0.7127 $\pm$ 0.0264
NdLinear 222 0.7906 $\pm$ 0.0142 0.7405 $\pm$ 0.0209
SST-2 Linear 1,544 0.8872 $\pm$ 0.0079 0.8867 $\pm$ 0.0080
NdLinear 222 0.8933 $\pm$ 0.0093 0.8932 $\pm$ 0.0073
D.2 Time Series
Time Series Forecasting. In our experiments using RNNs, we used a sequence length of 24 and a forecast horizon of 12 for all models. The models were trained for 100 epochs using the Adam optimizer with a learning rate of 0.02 and a batch size of 128. The dataset was split into training, validation, and evaluation sets with proportions of 60%, 20%, and 20%, respectively. We set the hidden size to 96, used a single recurrent layer, and applied a dropout rate of 0.3.
For Transformer-based Forecasting tasks, the experiments were conducted using a time series transformer model with a model dimension and hidden dimension both set to 32, a single transformer layer, and a dropout rate of 0.1. The GELU activation function was employed throughout, and the models were trained for 10 epochs with a batch size of 128 and a learning rate of 0.001. All models optimized using Adam and mean squared error as the loss function.
Time Series Classification. We set the hidden size of all RNN layers to 128 and used 3 recurrent layers, with a batch size of 32 and a learning rate of 0.005. Models were trained for 200 epochs using the Adam optimizer and cross-entropy loss.
D.3 Tabular Data
We compare the performance of the Linear and NdLinear models. The Linear Model uses two linear layers for feature extraction, while the NdLinear Model replaces them with NdLinear layers.
Classification. Target labels were one-hot encoded. The Linear Model utilized fully connected layers with input dimension $[11]$ and hidden dimension $[128]$, followed by ReLU and a final linear output layer. The NdLinear Model used custom NdLinear layers with input dimensions $[11,1]$ and hidden dimensions $[11,64]$, also followed by ReLU and a final linear output layer. Both models are trained over 40 epochs with a batch size of 32 and a learning rate of $0.0001$ using AdamW optimizer. Data was randomly shuffled and split into 80% training and 20% testing sets. Cross-entropy loss is used for training, and classification accuracy is used for model evaluation.
Regression. Target labels were kept as continuous values. The Linear Model utilized fully connected layers with input dimension $[14]$ and hidden dimension $[128]$, followed by ReLU and a final linear output layer. The NdLinear Model used custom NdLinear layers with input dimensions $[14,1]$ and hidden dimensions $[32,64]$, also followed by ReLU and a final linear output layer. Both models are trained over 40 epochs with a batch size of 32 and a learning rate of $0.0002$ using AdamW optimizer. Data was randomly shuffled and split into 80% training and 20% testing sets. MSE loss is used for both training and model evaluation.
Table 7: NdLinear with MLPs on tabular datasets. For classification (Cardio Disease), the metric is Accuracy (higher is better). For regression (Delivery Time), the metric is MSE (lower is better).
Dataset Task Method #Params Metric
Cardio Disease Classif. (Accuracy) Linear 18 306 0.7265
NdLinear 5 962 0.7321
Delivery Time Regress. (MSE) Linear 18 561 70.508
NdLinear 7 873 67.824
D.4 Vision
Image Classification with CNN. The NdLinear version uses three transforms with hidden dimensions of $(64,8,8)$, while the Linear version uses a single hidden dimension of $256$. Models were trained for 50 epochs using Adam optimizer (learning rate $0.001$), batch size 64, cross-entropy loss, and mixed-precision training on CUDA when available.
Table 8: Image classification with CNNs on CIFAR-10 (top-1 Acc.) and CIFAR-100 (top-5 Acc.). NdLinear achieves higher accuracy with fewer parameters.
Dataset #Params Method Accuracy
CIFAR-10 1.07M Linear 0.7426 $\pm$ 0.0025
65k NdLinear 0.7689 $\pm$ 0.0060
CIFAR-100 1.09M Linear 0.6587 $\pm$ 0.0075
433k NdLinear 0.7096 $\pm$ 0.0121
Vision Transformers (ViT) [Dosovitskiy et al., 2021]. Training used a batch size of 512, AdamW optimizer, learning rate $2.75\times 10^{-4}$ for 30 epochs, and a distillation temperature of 3. Input images ($224\times 224$) were augmented with random cropping and horizontal flipping.
Figure 2: NdLinear’s efficiency. Reduced ViT model parameter counts on CIFAR-10 and CIFAR-100 for a distillation task.
Table 9: NdViT vs. Naive ViT: Accuracy and parameter efficiency on CIFAR-10 (Acc@1) and CIFAR-100 (Acc@5). NdViT shows improved accuracy with fewer parameters.
Dataset Num. of Transformers NdViT (Ours) Naive
200 300 400 500
CIFAR10 3 Blocks 65.77 $\pm$ 0.47 67.53 $\pm$ 0.70 69.00 $\pm$ 1.27 62.09 $\pm$ 0.40
6 Blocks 68.48 $\pm$ 0.75 70.20 $\pm$ 0.73 72.03 $\pm$ 0.46 65.19 $\pm$ 0.64
9 Blocks 70.27 $\pm$ 0.35 71.50 $\pm$ 0.58 72.53 $\pm$ 0.54 68.52 $\pm$ 1.24
CIFAR100 3 Blocks 70.78 $\pm$ 1.36 73.10 $\pm$ 1.06 74.14 $\pm$ 1.66 69.34 $\pm$ 0.88
6 Blocks 73.60 $\pm$ 0.83 75.07 $\pm$ 0.14 76.37 $\pm$ 0.71 73.84 $\pm$ 0.39
9 Blocks 74.24 $\pm$ 0.32 75.52 $\pm$ 0.73 76.61 $\pm$ 0.26 75.60 $\pm$ 0.70
Diffusion Transformers (DiT) [Peebles and Xie, 2023]. We used a learning rate of $1\times 10^{-4}$ for $256\times 256$ images, mixed-precision (bfloat16) training, automatic gradient accumulation, and a batch size of 256, varying model depth and attention heads.
Table 10: FID-10k scores for DiT models: Pre-trained vs. NdLinear variants with fewer parameters.
NdLinear NdLinear Baseline
Parameter Count 619M 563M 674M
FID-10k 5.4876 5.9420 5.4109
Figure 3: DiT achieving lower (better) FID scores for image generation on ImageNet-100 when trained from scratch with comparable parameters.