### v3

# NdLinear: Preserving Multi-Dimensional Structure for Parameter-Efficient Neural Networks

Abstract
In deep learning, processing multidimensional inputs (e.g., images, medical scans, and time series) is an important task that often requires flattening the inputs. We introduce NdLinear, a drop-in replacement for linear layers that operates directly on tensors, requiring no flattening. By applying transformations separately along each dimension, NdLinear preserves native data structure while achieving dramatic parameter reductions, often by orders of magnitude, with minimal memory overhead. We prove NdLinear maintains expressivity through structured Tucker decomposition while preserving VC-dimension scaling. Extensive experiments demonstrate NdLinear’s capacity to achieve significant parameter reductions with substantial wall-clock efficiency gains and minimal memory overhead. For instance, our NdLinear-LoRA matches or exceeds standard LoRA on language reasoning tasks using up to $9\times$ fewer parameters. Experiments across CNNs, RNNs, Transformers, and MLPs on vision, language, time-series, and tabular tasks consistently demonstrate NdLinear’s efficiency gains. While excelling at axis-separable tasks, NdLinear has limitations with entangled spatial interactions. By processing data in its original N-dimensional form, NdLinear provides a theoretically grounded, practical component for building more efficient neural architectures.
footnotetext: NdLinear is licensed by Ensemble AI under the Apache 2.0 license. Our technical contributions and research findings are free for use by the open-source community.
1 Introduction
Deep learning excels at processing multidimensional data, such as medical scans, videos and sensor arrays. Yet, a fundamental inefficiency persists throughout modern architectures. Consider what happens when a 32×32×32 voxel tensor reaches a linear layer: it gets flattened into a 32,768-dimensional vector, scrambling the careful spatial organization that earlier layers worked to extract. This pattern repeats everywhere linear layers appear: CNN classification heads flatten spatial feature maps, Transformer MLPs flatten structured token representations, and RNN projections flatten temporal features. Beyond this massive inefficiency, flattening forces networks to waste capacity rediscovering that adjacent voxels are related, that temporal continuity matters, and that feature dimensions have distinct meanings. While convolutions preserve local structure, they cannot replace linear layers for global reasoning, channel mixing, or final predictions. This leaves a critical gap: no existing layer can perform general linear transformations while respecting multidimensional structure.
To address this, we introduce NdLinear, a novel linear layer that operates natively on N-D tensors. Unlike vanilla linear layers, NdLinear processes data in its original N-D form by applying distinct linear transformations sequentially along each tensor dimension. This dimension-wise approach inherently preserves the data’s N-D structure.
Preserving data’s native organization allows NdLinear to offer compelling advantages: It enhances representational power by maintaining structural integrity and facilitating a more natural information flow. It dramatically reduces parameter counts (often by orders of magnitude vs. flattened layers) and ensures efficient computation without sacrificing performance. These efficiencies lead to faster training/inference and lower memory usage. As a general-purpose layer, NdLinear is a versatile, drop-in replacement for vanilla linear layers in many architectures, benefiting diverse domains by enabling models to better leverage N-D data characteristics.
Our work makes the following primary contributions, substantiated by extensive empirical validation:
  • We articulate the critical limitations of flattening in conventional linear layers for N-D data.
  • We introduce and formulate NdLinear, a novel N-D linear layer using sequential, dimension-wise transformations to preserve data structure and achieve substantial parameter efficiency.
  • We empirically demonstrate through comprehensive evaluations that NdLinear significantly enhances model performance, or matches it with drastically fewer parameters, when replacing vanilla linear layers across diverse architectures (CNNs, RNNs, Transformers, MLPs) and data domains, including vision, language, time-series, and tabular data.
  • We analyze the significant parameter and computational efficiency gains originating from NdLinear’s dimension-wise transformation principle.
By demonstrating that preserving N-D structure yields both theoretical and practical advantages, NdLinear challenges the conventional wisdom that flattening is a necessary evil in neural architectures.
2 Related Work
Neural networks process multidimensional data through three main approaches, each with fundamental limitations:
Flattening-based Methods.
Standard linear layers reshape N-D tensors to 2D matrices, destroying spatial and dimensional relationships. This forces networks to relearn structure from scrambled features, requiring excessive parameters. For example, a $32\times 32\times 32$ tensor needs $\sim 10^{9}$ parameters when flattened, versus $\sim 10^{3}$ if structure were preserved.
Specialized Architectures.
Convolution layers excel at 2D/3D spatial patterns but become unwieldy beyond three dimensions. Depthwise separable convolutions (Chollet, 2017a) and axial attention (Ho et al., 2019) reduce complexity through factorization but remain tied to specific spatial operations rather than general tensor transformations.
Tensor Decomposition Layers.
Recent work applies tensor algebra to neural networks. Tensor Contraction Layers (TCL) (Kossaifi et al., 2020) use multilinear maps to compress activations to lower dimensions. Tensor Regression Layers (TRL) (Kossaifi et al., 2020) parameterize predictions using Tucker or Tensor-Train formats. Tensor-Train layers (Novikov et al., 2015) decompose weights for compression. While efficient, these methods focus on specific tasks, dimensionality reduction or regression, rather than general-purpose transformation.
Gap in Current Methods.
Despite extensive work on structured layers, no method provides a true drop-in replacement for linear layers that: (1) operates directly on N-D tensors without flattening, (2) supports flexible dimension-wise transformation (expansion or compression), (3) integrates seamlessly with modern architectures (bias, normalization, dropout). This gap motivates NdLinear, which we introduce in Section 3.
We provide detailed mathematical comparisons between NdLinear and existing tensor methods (TCL, TRL, $T$ decomposition) in Appendix A, demonstrating how our approach differs fundamentally in design philosophy and implementation, leading to the superior empirical results shown in Section 4.
3 Linear Transformation Preserving N-dimensional Information
Vanilla linear layers cannot process input tensors directly. They require transforming the inputs into 2D matrices, destroying the original N-D structure. For an N-D input tensor $X\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$ with batch size $B$ and feature dimensions $(D_{1},\ldots,D_{N})$, standard layers flatten it to $\mathbb{R}^{B\times\prod_{i}D_{i}}$ before applying a linear transformation. NdLinear suggests an alternative approach that transforms $X$ directly to $Y\in\mathbb{R}^{B\times H_{1}\times\cdots\times H_{N}}$ without flattening, where each dimension $D_{k}$ maps to $H_{k}$, while preserving the N-D structure throughout the transformation.
The NdLinear transformation processes N-D tensors by applying separate linear transformations along each of their feature dimensions sequentially. This contrasts with vanilla linear layers, which flatten the N-D tensor into a 2D matrix, losing the original multidimensional structure. By operating on each dimension independently while preserving the others, NdLinear retains the inherent structural information of the data throughout the transformation.
Conceptually, for an input tensor $X\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$, NdLinear learns $N$ distinct weight matrices $W_{1},\ldots,W_{N}$, where each $W_{k}\in\mathbb{R}^{D_{k}\times H_{k}}$ transforms the $k$-th dimension of the input from its original size $D_{k}$ to a new size $H_{k}$, with optional bias vectors $b_{k}\in\mathbb{R}^{H_{k}}$ per dimension. The transformation is applied iteratively: the output from transforming dimension $k$ becomes the input for transforming dimension $k+1$. The procedure is detailed in Algorithm 1.
Algorithm 1 NdLinear Transformation
0: Input tensor $X\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$ (batch size $B$, original feature dimensions $D_{1},\ldots,D_{N}$), Target output feature dimensions $H_{1},\ldots,H_{N}$, Learnable weight matrices $W_{k}\in\mathbb{R}^{D_{k}\times H_{k}}$ for $k=1,\ldots,N$, Optional learnable bias vectors $b_{k}\in\mathbb{R}^{H_{k}}$ for $k=1,\ldots,N$.
 Let $X_{\text{out}}\leftarrow X$.
for $k=1$ to NN do
  Let current shape of $X_{\text{out}}$ be $(B,S_{1},\ldots,S_{N})$, where $S_{j}=H_{j}$ for $j<k$, and $S_{j}=D_{j}$ else.
  The $k$-th feature dimension (original size $D_{k}$, current size $S_{k}=D_{k}$) is targeted for transformation.
  Permute $X_{\text{out}}$ to move its $k$-th feature dimension to the last position. Shape becomes $(B,S_{1},\ldots,S_{k-1},S_{k+1},\ldots,S_{N},D_{k})$.
  Reshape $X_{\text{out}}$ to a 2D matrix $X_{\text{matrix}}$ of shape $\left(B\cdot\prod_{j\neq k}S_{j},D_{k}\right)$.
  Apply linear transformation: $Y_{\text{matrix}}\leftarrow X_{\text{matrix}}W_{k}+b_{k}$. ($Y_{\text{matrix}}$ has shape $\left(B\cdot\prod_{j\neq k}S_{j},H_{k}\right)$).
  Reshape $Y_{\text{matrix}}$ back to N-D form: $(B,S_{1},\ldots,S_{k-1},S_{k+1},\ldots,S_{N},H_{k})$.
  Permute dimensions to place $H_{k}$ (the new size of the $k$-th feature dimension) back into the $k$-th feature position. Shape becomes $(B,S_{1},\ldots,S_{k-1},H_{k},S_{k+1},\ldots,S_{N})$.
  Update $X_{\text{out}}$ with the result of this transformation.
end for
return $X_{\text{out}}$, now of shape $\mathbb{R}^{B\times H_{1}\times\cdots\times H_{N}}$.
In practice, these operations (transposing, reshaping, matrix multiplication, then inverse reshaping and transposing) can be efficiently implemented using standard tensor library functions like torch.tensordot or einsum. The key is that each weight matrix $W_{k}$ only transforms dimension $D_{k}$ to $H_{k}$. This operation modifies all entries along the $k$-th mode, performing the same linear transformation on each mode-$k$ fiber of the tensor.
This sequential application can be expressed using tensor notation as a series of mode-$k$ tensor-matrix products (Kolda and Bader, 2009): $Y=X\times_{1}W_{1}\times_{2}W_{2}\cdots\times_{N}W_{N}$ , where each product $X\times_{k}W_{k}$ transforms the $k$-th mode of the tensor using matrix $W_{k}$. (Biases $b_{k}$ are added after each mode-$k$ product). The intermediate result of $X\times_{k}W_{k}$ becomes the input for the $\times_{k+1}W_{k+1}$ product.
3.1 Preserving The Expressiveness of Vanilla Linear Layers
The dramatic parameter reduction of NdLinear raises a natural question: does this efficiency sacrifice model expressivity? We show that despite using fewer parameters, NdLinear maintains sufficient representational capacity.
Theoretical Guarantee. We analyze expressivity through VC-dimension, which measures a model’s capacity to fit arbitrary patterns. Following Bartlett et al. (2019), networks with $P$ parameters have VC-dimension $\Theta(P\log P)$.
Theorem 3.1 (Informal; see Appendix C for formal statement).
An NdLinear network with $P_{\text{nd}}=d(a+b+c)$ parameters for tensor dimensions $(a,b,c)$ and hidden dimension $d$ maintains VC-dimension $\Theta(P_{\text{nd}}\log P_{\text{nd}})$ as $d\to\infty$, matching the scaling of vanilla linear layers with $P_{\text{std}}$ parameters.
Practical Implication. While NdLinear uses fewer parameters, the reduction is polynomial and not exponential in the VC-dimension bound. This theoretical guarantee is validated empirically: NdLinear often improves performance despite parameter reduction, suggesting the structured factorization acts as beneficial regularization rather than a limiting constraint.
Empirical Evidence: Representation Compression. To understand why fewer parameters improve performance, we measured von Neumann entropy of learned representations. On the Radius Bump task F.3 across varying difficulties, NdLinear consistently produces 15-30% lower entropy than parameter-matched dense networks while achieving equal or better test MSE (Appendix F.3). This lower entropy, indicating more compressed representations, aligns with recent theoretical work showing compressed representations generalize better (Skean et al., 2025). NdLinear’s structured factorization inherently eliminates redundancy while preserving task-relevant information, explaining why dramatic parameter reduction enhances rather than hurts performance.
3.2 Fewer Learnable Parameters And Wall-Clock Speedups
A key advantage of NdLinear is its parameter efficiency. Consider transforming $X\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$ to $Y\in\mathbb{R}^{B\times H_{1}\times\cdots\times H_{N}}$.
Parameter Reduction. Vanilla linear layers require $(\prod_{i}D_{i})\times(\prod_{i}H_{i})$ parameters, which grow exponentially with dimensionality. NdLinear requires only $\sum_{i=1}^{N}(D_{i}H_{i})$ parameters, which grow linearly. For example, transforming a $32\times 32\times 32$ tensor to the same size requires $\sim 10^{9}$ parameters for vanilla linear but only $3,072$ for NdLinear, a reduction of six orders of magnitude.
Computational Complexity. NdLinear’s FLOPs scale as $\mathcal{O}(B\cdot N\cdot D^{N+1})$ for cubic tensors with $D_{i}=H_{i}=D$, compared to $\mathcal{O}(B\cdot D^{2N})$ for vanilla linear layers. This yields order-of-magnitude speedups that increase with tensor dimensionality. The exact FLOP count is:
$\text{FLOPs}_{\text{NdLinear}}=B\sum_{k=1}^{N}\left(\prod_{j<k}H_{j}\cdot\prod_{j>k}D_{j}\cdot D_{k}H_{k}\right).$
Memory Efficiency. While theoretical analysis bounds peak memory overhead at $1/N$ of baseline, which is typically $\leq 33\%$ for 3D tensors (proof in Appendix C.2), empirical measurements show much smaller overhead in practice: only 1.1-2.0% increased peak memory across diverse architectures, with training time overhead below 1.6% (Section 4.5).
3.3 Inductive Bias and Domain Alignment
NdLinear’s efficiency stems from a strong inductive bias: it assumes dimensions can be transformed independently. This bias has clear benefits and limitations.
The Separability Assumption. NdLinear implements a rank-1 Tucker decomposition, assuming the transformation can be factorized across dimensions. Crucially, this structure persists through ReLU-like activations. For 1-homogeneous activations $\sigma$ (ReLU, GELU):
$\sigma(T(x))=\sum_{r=1}^{R}\bigotimes_{k=1}^{m}\sigma(u_{r}^{(k)}),$
where output rank remains bounded even after nonlinearities. This ensures the separability bias persists through network depth rather than degrading.
Domain Alignment. NdLinear excels when data dimensions represent independent factors, such as spectrograms (frequency × time), sensor arrays (sensor × time), and tabular data, where features have separable effects. However, it struggles with dense cross-dimensional interactions like XOR patterns, checkerboards, or highly entangled spatial features where dimensions cannot be meaningfully separated.
3.4 Quantifying the Trade-off: Dial-a-Bias Experiment
To precisely characterize when NdLinear’s bias helps versus hurts, we designed a controlled experiment interpolating between separable and entangled patterns.
We create synthetic targets that blend separable and entangled components:
$y=(1-\alpha)\cdot f_{\text{separable}}(X)+\alpha\cdot f_{\text{entangled}}(X),\quad\alpha\in[0,1],$
where $X\in[0,1]^{32\times 32}$ has i.i.d. uniform entries. The separable component $f_{\text{separable}}$ aggregates per-axis statistics (row/column means, variances), while $f_{\text{entangled}}$ requires cross-axis interactions (XOR patterns, checkerboards).
Figure 1: NdLinear excels on separable tasks but struggles with entangled patterns Performance comparison as task structure varies from purely separable ($\alpha=0$) to fully entangled ($\alpha=1$). The crossover at $\alpha\approx 0.45$ indicates NdLinear outperforms dense MLPs when tasks have <<45% entanglement. This provides clear deployment guidance: use NdLinear for axis-aligned domains (spectrograms, time series, tabular data) but prefer dense layers for spatially entangled tasks (dense vision, XOR-like patterns).
Experimental Protocol. We trained both NdLinear and dense MLPs (2 hidden layers, 256 units) on 10,000 samples for each $\alpha\in\{0,0.1,...,1.0\}$.
Results. Figure 1 reveals a sharp crossover at $\alpha\approx 0.45$. NdLinear achieves 49× lower MSE for purely separable tasks ($\alpha=0$) but 64× worse for fully entangled tasks ($\alpha=1$). This provides clear guidance on the use of NdLinear, where the target structure aligns with axis-separable patterns.
4 Experimental Results
We evaluate NdLinear as a drop-in replacement for vanilla linear layers across language, vision, time-series, and tabular domains. Our experiments span CNNs, RNNs, Transformers (including LLMs), MLPs, and DiTs, with model scales from 65K to 8B parameters on over 20 datasets. We also compare NdLinear against alternative structured layers including Tensor Regression Layers (TRL/TCL) and Tensor-Train (TT) decompositions (Novikov et al., 2015; Kossaifi et al., 2020)
Universal findings: NdLinear reduces parameters by 50-95% while maintaining or improving performance across all tested configurations. When compared to other structured approaches, NdLinear consistently achieves superior accuracy with lower computational overhead. We present domain-specific results below, with ablations in Section  4.5 and full details in Appendix E.
4.1 Natural Language Processing
4.1.1 Parameter-Efficient Finetuning with NdLinear-LoRA
We introduce NdLinear-LoRA, replacing LoRA’s low-rank matrices $A,B$ with a single NdLinear adapter. Concretely, the LoRA update $\Delta W=BA$ is swapped for an NdLinear module that applies sequential mode-wise transforms on the reshaped activation tensor; all other LoRA mechanics (zero-init, scaling $\alpha$, residual addition, and merge-at-inference) remain unchanged. This preserves the native $N$-D structure during adaptation while keeping the same drop-in interface and training pipeline as standard LoRA.
We fine-tuned Qwen3-1.7B (Yang et al., 2025) and LLaMA3-8B (Dubey et al., 2024) on mathematical reasoning (OpenMathInstruct-1 (Toshniwal et al., 2024)) and commonsense QA (CommonsenseQA training split (Talmor et al., 2019)), then evaluated on GSM8K, MultiArith, CSQA, ARC-Easy, ARC-Challenge, and BoolQ..
Results. Table 1 shows NdLinear-LoRA achieves superior performance with dramatically fewer parameters. On Qwen3-1.7B with 87% fewer parameters (1.15M vs 8.72M), it improves GSM8K by 11.9 points and MultiArith by 7.8 points over standard LoRA. On LLaMA3-8B with 9× fewer parameters (2.26M vs 20.97M), it achieves the best CSQA (82.9%) and ARC-Challenge (76.6%) scores. This suggests structured transformations capture task-specific patterns more efficiently than low-rank factorization.
Table 1: NdLinear-LoRA demonstrates significant parameter efficiency for LLM finetuning, achieving comparable or improved accuracy over standard LoRA methods with up to $9\times$ fewer trainable parameters. Accuracy ($\uparrow$) after LoRA of Qwen3-1.7B and LLaMA3-8B models.
Math CS Reasoning
Model Method Params GSM8K M.Arith CSQA ARC‑e ARC‑c BoolQ
Qwen3‑1.7B LoRA (r=4r{=}4) 4.36M 45.6 88.9 80.4 91.9 79.4 79.7
LoRA (r=8r{=}8) 8.72M 40.3 82.2 80.9 91.8 79.3 80.8
NdLinear‑LoRA 1.15M 52.2 90.0 81.0 92.2 78.3 79.7
LLaMA3‑8B LoRA (r=4r{=}4) 10.48M 50.5 84.4 80.6 90.4 76.3 85.1
LoRA (r=8r{=}8) 20.97M 51.6 81.1 81.7 89.0 73.6 76.5
NdLinear‑LoRA 2.26M 40.2 80.0 82.9 90.9 76.6 80.5
4.1.2 Language Model Pretraining
We pretrained OPT models (Zhang et al., 2022) (124M and 350M parameters) from scratch on BookCorpus and Wiki40B-English, replacing feedforward linear layers with NdLinear. Despite fewer parameters, NdLinear variants achieve lower perplexity (Table 5 in the Appendix). The improvement scales with model size: perplexity gap increases from 0.215 (OPT-Small) to 0.361 (OPT-Mid), suggesting NdLinear’s benefits grow with scale.
4.1.3 Text Classification
We evaluated NdLinear in BERT’s classification head (Devlin et al., 2019) on SST-2 sentiment analysis (Socher et al., 2013) and CoLA grammatical acceptability (Warstadt et al., 2018) tasks. We replaced BERT’s standard two-layer classification head with a single NdLinear layer followed by a final linear projection.
Results. As shown in Table 7 in Appendix E, NdLinear improves both accuracy and ROC-AUC on both datasets while reducing the classification head parameters by approximately 85%. On SST-2, accuracy improves from 88.72% to 89.33%; on CoLA, from 77.90% to 79.06%.
4.2 Time Series Analysis
Multivariate Time Series Forecasting. We evaluated NdLinear in RNNs and Transformers for 12-hour ahead forecasting on four ETT (Electricity Transformer Temperature) datasets (Zhou et al., 2021), using 24 hours of historical input. For RNNs, we replaced vanilla linear layers with NdLinear. For Transformers, we replaced linear layers in feedforward blocks with NdLinear.
Results. Table 3 shows NdLinear consistently improves forecasting accuracy across all datasets while using 50% fewer parameters. For RNNs on ETTh1, MSE decreases from 0.290 to 0.088 (70% reduction). For Transformers on ETTh2, MSE decreases from 0.0226 to 0.0158 (30% reduction). These improvements demonstrate NdLinear’s natural alignment with temporal-feature decomposition in multivariate time series.
Time Series Classification. We applied Transformer models with NdLinear to six UCR time series classification datasets (Dau et al., 2019), replacing vanilla linear layers in transformer blocks with NdLinear layers.
Table 2: NdLinear significantly enhances Transformer-based time series classification, achieving superior F1 scores across UCR datasets with comparable or substantially fewer parameters than baseline models. F1 Score and efficiency (# Params) of NdLinear compared to a standard Base Model and a parameter-comparable Base Model (whose parameters are aligned with NdLinear’s).
Model Metric ECGFiveDay HeartBeat Chlorine Conc. ECG5000 LSST Sleep
Base Model Params 3 363 4 323 12 900 12 966 29 343 12 966
F1 Score 0.7668 0.7250 0.4436 0.8886 0.5928 0.4911
Base Model Params 1 779 2 739 6 660 6 726 15 375 6 726
F1 Score 0.7624 0.7214 0.4440 0.8878 0.5907 0.4897
NdLinear Params 1 804 2 752 6 709 6 823 15 472 6 823
F1 Score 0.7783 0.7363 0.4571 0.9058 0.6486 0.4978
Table 3: NdLinear substantially improves time-series forecasting accuracy and dramatically reduces parameter counts in both RNN and Transformer models on Electricity Transformer Temperature (ETT) datasets. With NdLinear, model parameters are reduced by 50%, alongside consistent gains in accuracy. Metrics reported are mean $\pm$ standard deviation over three experimental runs.
RNN Transformer
Dataset Method (Params: RNN / Trans.) MSE MAE MSE MAE
ETTh1 Linear (20.5k / 138k) 0.2900 $\pm$ 0.0170 0.4060 $\pm$ 0.0246 0.0217 $\pm$ 0.0001 0.1158 $\pm$ 0.0004
NdLinear (9.6k / 70k) 0.0880 $\pm$ 0.0115 0.2204 $\pm$ 0.0105 0.0173 $\pm$ 0.0003 0.0995 $\pm$ 0.0005
ETTh2 Linear (20.5k / 138k) 0.2636 $\pm$ 0.0949 0.3955 $\pm$ 0.0748 0.0226 $\pm$ 0.0001 0.1229 $\pm$ 0.0003
NdLinear (9.6k / 70k) 0.1536 $\pm$ 0.0137 0.2831 $\pm$ 0.0119 0.0158 $\pm$ 0.0019 0.0995 $\pm$ 0.0071
ETTm1 Linear (20.5k / 138k) 0.0187 $\pm$ 0.0012 0.0926 $\pm$ 0.0039 0.0180 $\pm$ 0.0001 0.1005 $\pm$ 0.0005
NdLinear (9.6k / 70k) 0.0174 $\pm$ 0.0017 0.0894 $\pm$ 0.0039 0.0161 $\pm$ 0.0006 0.0936 $\pm$ 0.0027
ETTm2 Linear (20.5k / 138k) 0.0148 $\pm$ 0.0007 0.0825 $\pm$ 0.0047 0.0151 $\pm$ 0.0000 0.0965 $\pm$ 0.0001
NdLinear (9.6k / 70k) 0.0139 $\pm$ 0.0009 0.0797 $\pm$ 0.0039 0.0141 $\pm$ 0.0001 0.0929 $\pm$ 0.0004
Results. Table 8 in Appendix E shows NdLinear reduces parameters by up to 47% while consistently improving F1 scores across all datasets. For example, on LSST, F1 improves from 0.5928 to 0.6486 with 47% fewer parameters. Even compared to parameter-matched baselines (Base Model*), NdLinear variants achieve superior performance, highlighting that the architectural advantage extends beyond mere parameter reduction.
4.3 Tabular Data
We evaluated NdLinear in MLPs on tabular classification (Cardio Disease dataset) (Sulianova, 2025) and regression (Food Delivery Time dataset) (Kumar, 2025). We replaced vanilla linear layers in two-layer feature extractors with NdLinear layers.
Results. Table 2 in shows NdLinear improves performance while dramatically reducing parameters. For Cardio Disease classification, NdLinear achieves higher accuracy (73.21% vs 72.65%) with 67% fewer parameters. For Food Delivery Time regression, NdLinear reduces MSE from 70.51 to 67.82 with 58% fewer parameters. These results demonstrate that NdLinear’s dimensional factorization aligns well with tabular data where features often have independent effects.
4.4 Computer Vision
4.4.1 Vision Transformer Distillation
We created student ViTs (NdViT) by replacing all linear layers in feedforward blocks with NdLinear, distilling from a pretrained ViT-B/16 teacher (Dosovitskiy et al., 2021). We tested various embedding dimensions (200, 300, 400) and depths (3, 6, 9 blocks) on CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009).
Results. Table 10 and Figure 3 in Appendix E show NdViT consistently outperforms standard ViT students while using 26-68% fewer parameters. Notably, NdViT with embedding dimension 200 surpasses a standard ViT with dimension 500, demonstrating superior parameter efficiency. The performance gap increases with model depth, suggesting NdLinear’s benefits compound in deeper architectures.
4.4.2 CNN Image Classification
We replaced the penultimate linear layer in CNNs with NdLinear on CIFAR-10 and CIFAR-100, and compared against Tensor Regression Layers (TRL/TCL) and Tensor-Train (TT) layers.
Results. NdLinear achieves higher accuracy with significantly fewer parameters on both datasets (Table 9 in Appendix E). On CIFAR-10, NdLinear improves accuracy by 2.6% while reducing parameters by 94%. On CIFAR-100, it improves accuracy by 5.1% with 60% fewer parameters.
Comparison with Structured Methods. Table 4 compares NdLinear against TRL/TCL and TT layers on CIFAR-100. NdLinear achieves superior accuracy (71.3%) compared to TRL/TCL (69.4%) and TT (56.2%), while using fewer parameters (434K vs 548K vs 769K), lower FLOPs (0.84G vs 3.97G vs 5.25G), 6× lower latency than TT (0.98ms vs 5.87ms), and minimal GPU memory overhead (35.16 MB vs 35.60 MB vs 100.44 MB). This demonstrates NdLinear’s practical advantages over alternative structured approaches.
Table 4: NdLinear outperforms existing structured tensor methods. Comparison with Tensor Regression Layers (TRL/TCL) and Tensor-Train decomposition on CIFAR-100 classification, replacing the CNN’s penultimate linear layer.
Method Mem (MB) Acc@5 Latency (s) FLOPs (G) Params
NdLinear 35.16 0.7133 0.000976 0.843 433,588
TRL/TCL 35.60 0.6935 0.001116 3.97 548,032
$T$ 100.44 0.5617 0.005871 5.25 769,316
4.4.3 Generative Modeling with Diffusion Transformers (DiT)
We evaluated NdLinear in DiTs through two approaches: training from scratch on ImageNet-100 (Russakovsky et al., 2015) and modifying pre-trained DiT models from (Jin and Xie, 2024). For training from scratch, we replaced linear layers in the timestep embedder. For pretrained models, we replaced linear layers in both timestep embedders and attention MLPs.
Results. When trained from scratch, NdLinear variants achieve lower (better) FID scores with comparable parameters (Figure 4 in Appendix E). For pretrained models, NdLinear reduces parameters from 674M to 619M while maintaining similar FID scores. demonstrating that parameter reduction doesn’t sacrifice generation quality.
Modifying Linear Layers in Pre-trained DiT Models
To evaluate performance maintenance with parameter reduction, we replaced standard Linear layers with NdLinear in pre-trained DiTs. This was done in both the timestep embedder and the MLP components of attention heads.
Results. Table 11 in Appendix E shows that NdLinear-based models (619M and 563M) achieve FID scores comparable to the larger baseline 674M model, despite significant parameter reductions.
4.5 Ablation Studies
We conducted extensive ablations to understand NdLinear’s design choices and practical considerations (full details in Appendix F).
4.5.1 Design Choices
Per-mode Bias Terms. Including bias terms for each dimension transformation significantly improves performance, with benefits increasing at larger widths. On the Radius Bump task, bias terms improve MSE by 4.5% at width 32 and 15.2% at width 128, with negligible parameter overhead (<5%) (Table F.1 in the Appendix). This suggests per-mode biases help capture dimension-specific offsets. More information on the taks can be found in Appendix F.3
Axis Ordering Robustness. NdLinear shows remarkable robustness to transformation order. Permuting axes (original, reverse, random) cause only 4% accuracy variation on CIFAR-100, with random ordering achieving 96% of baseline performance (Table F.2 in the Appendix). This suggests NdLinear learns relatively axis-independent features, simplifying deployment.
Hidden dimensions offer a direct trade-off between efficiency and expressivity. Doubling hidden size improved accuracy by 2% while still using 40% fewer parameters than dense layers.
4.5.2 Practical Considerations
Hyperparameter Robustness. NdLinear requires no special tuning. Across learning rates (0.001, 0.01), batch sizes (32, 64, 128), and hidden configurations, accuracy varies only by 11 percentage points (64-75%) on CIAFR-10 (Table F.4 in the Appendix). Even the worst configuration (LR=0.01, batch=32) achieves 64% accuracy with 6× fewer parameters than dense layers, demonstrating robustness to suboptimal settings.
Computational Overhead. Despite sequential processing, measured overheads are minimal: peak memory increases by 1.1-2.0% (CIFAR CNN: 35.17→36.91 MB), training time by 0.6-1.6% (47.2→47.8s per epoch). Inference latency remains comparable due to reduced FLOPs offsetting sequential operations (Table F.6 in the Appendix).
Sample Efficiency. The separability bias dramatically affects data efficiency (Section F.5). On our synthetic tasks with $\alpha=0.1$ separability, NdLinear reaches target error with 5× fewer samples (2,000 vs 10,000). Conversely, at $\alpha=0.9$ entanglement, it requires 1.7× more samples (25,000 vs 15,000), quantifying the bias-variance trade-off. The representation analysis in Section 3.1 provides additional mechanistic insight into these efficiency gains.
Architectural Comparisons. As shown in Section 4.4.2, NdLinear outperforms alternative structured layers (TRL/TCL, TT) on all metrics, validating our design choices.
Key Takeaway.
NdLinear is robust to implementation details and requires no special tuning, where standard hyperparameters work well. The critical decision is architectural: whether the task exhibits sufficient separability to benefit from NdLinear’s bias.
5 Conclusion
We introduced NdLinear, a linear layer that operates directly on N-dimensional tensors through sequential dimension-wise transformations. This simple change, by processing tensors in their native form rather than flattening them, yields dramatic improvements across modern deep learning.
Our extensive experiments demonstrate that NdLinear reduces parameters by 50-95% while maintaining or improving performance across all tested configurations. The results reveal striking consistency: from NdLinear-LoRA fine-tuning LLaMA-8B with 9× fewer parameters while improving reasoning accuracy, to achieving 70% lower error in time-series forecasting with half the parameters, to outperforming alternative structured methods (TRL/TCL, TT). This universality, spanning language, vision, time-series, and tabular domains, validates dimension-wise factorization as a fundamental principle.
Theoretically, we proved NdLinear maintains expressivity through preserved VC-dimension scaling, explaining why dramatic parameter reduction doesn’t sacrifice performance. Our controlled experiments precisely quantified when the method’s inductive bias helps versus hurts: NdLinear excels when data exhibits axis-separable structure, but struggles with highly entangled patterns. This theoretical understanding, combined with our ablations showing robustness to hyperparameters and negligible computational overhead (<2%), provides clear deployment guidance.
Beyond its immediate practical benefits, NdLinear challenges a fundamental assumption in neural architecture design: that flattening is necessary for linear transformations. By demonstrating that structure-preserving operations consistently outperform their flattened counterparts, we reveal that the ubiquitous practice of tensor flattening has been systematically discarding valuable inductive biases. As neural networks evolve to process increasingly complex multidimensional data, from volumetric medical imaging to spatiotemporal climate models, NdLinear’s principle of native dimensional processing offers not just an optimization, but a paradigm shift toward architectures that inherently respect and leverage the structure of our multidimensional world.
Appendix A Detailed Comparison with Related Methods
NdLinear’s design as a structure-preserving, parameter-efficient linear layer for N-D tensors is informed by, yet distinct from, several established concepts in machine learning and tensor algebra. We detail these relationships below.
A.1 Tensor Decomposition Methods
NdLinear’s core mechanism employs mode-wise tensor-matrix products prominent in Tucker decomposition [Tucker, 1966, Kolda and Bader, 2009]. However, unlike classical tensor decomposition methods that primarily analyze or compress static data or pre-existing weight tensors [Novikov et al., 2015, Newman et al., 2024], NdLinear integrates these mode-wise operations as a learnable, dynamic layer within a network. Its purpose is efficient forward transformation of activations while preserving N-D structure, not data analysis or post-hoc model compression.
NdLinear can be viewed as a hand-crafted factorization of a fully-connected weight matrix. Specifically, the full weight matrix $W_{\text{full}}$ implicitly has a Kronecker product structure derived from mode-wise matrices $\{W_{1},\ldots,W_{n}\}$. This corresponds to a rank-1 Tucker decomposition without a core tensor (or equivalently, a core of rank 1 in each mode).
The main trade-off is expressiveness vs. efficiency. NdLinear’s decomposition is low-rank in a multilinear sense—it cannot represent arbitrary non-factorizable interactions between dimensions. More flexible decompositions (full Tucker or higher-rank tensor decompositions) capture more interactions but require significantly more parameters compared to NdLinear’s simple sum $\sum_{i}D_{i}H_{i}$ and can be more challenging to train, sometimes needing special initialization or multi-stage training.
If needed, one can extend NdLinear by increasing the factorization rank (e.g., learning multiple $W_{i}$ matrices per mode and summing their effects, analogous to a rank-$R$ core). However, our experiments show the simple rank-1 version already performs well. Each $W_{i}$ clearly indicates how dimension $i$ is transformed, providing better interpretability than general tensor decomposition methods that disperse transformations across multiple factors.
A.2 Factorized Convolutions and Axial Operations
Neural nets have long exploited axis-wise structure to cut parameters and impose useful priors. In CNNs, an image tensor $X\!\in\!\mathbb{R}^{H\times W\times C}$ is not flattened; instead, convolutions slide local kernels over the two spatial axes $(H,W)$ while mixing channels $C$. In sequence models, a tensor $X\!\in\!\mathbb{R}^{L\times C}$ (length $L$, features $C$) is processed along the sequence axis $L$ (e.g., attention/conv) while mixing feature channels $C$ within positions. Below we recap common factorized operators and their limitations.
Definitions.
Channels are feature maps carried alongside a position-like axis (e.g., RGB planes or intermediate filters in CNNs; hidden features per token in sequences). Sequence (or spatial) dimensions are position-like axes along which locality or ordering matters (e.g., time $L$, image height $H$, width $W$).
Grouped convolutions.
Grouped convs restrict each filter to operate on a subset of channels, partitioning $C$ into groups to reduce parameters and FLOPs [Krizhevsky et al., 2012, Xie et al., 2017]. Limitation: channel mixing is constrained within groups; cross-group interactions require additional layers, and the grouping choice is a manual architectural prior.
Depthwise separable convolutions.
Depthwise separable convs factor a full conv into a depthwise spatial conv per channel followed by a $1{\times}1$ (pointwise) conv that mixes channels [Chollet, 2017b, Howard et al., 2017]. This yields large FLOP/parameter savings with strong accuracy. Limitations: the factorization is tied to 2D spatial structure; expressivity hinges on the pointwise mixer; extending beyond standard spatial axes typically needs custom kernels.
Axial (factorized) attention.
Axial attention decomposes 2D/3D attention into a sequence of 1D attentions applied along one axis at a time (e.g., height then width), cutting quadratic costs while preserving long-range interactions along each axis [Ho et al., 2019, Wang et al., 2020, Yan et al., 2023]. Limitations: axis order becomes an architectural prior; full joint interactions across axes emerge only after stacking, and costs can still be high for very long axes.
CNNs.
CNNs are the canonical instance of axis-aware processing: they exploit spatial locality (factorization in $(H,W)$) and defer heavy channel mixing to $1{\times}1$ pointwise layers. Grouped/depthwise variants intensify this factorization to further reduce compute.
NdLinear in this landscape.
NdLinear generalizes the factorization principle beyond conv/attention mechanics: given an $n$-D activation $X\!\in\!\mathbb{R}^{D_{1}\times\cdots\times D_{n}}$, it systematically factorizes a linear map into mode-wise transforms $D_{i}\!\to\!H_{i}$ and applies them sequentially, producing a structured output $Y\!\in\!\mathbb{R}^{H_{1}\times\cdots\times H_{n}}$. Thus, instead of a single dense matrix $\mathbb{R}^{\prod_{i}D_{i}}\!\to\!\mathbb{R}^{\prod_{i}H_{i}}$, NdLinear uses $n$ small matrices with $\sum_{i}D_{i}H_{i}$ parameters, preserving the N-D shape throughout. Unlike grouped/depthwise convs (tied to spatial kernels) or axial attention (tied to attention mechanics and axis ordering), NdLinear: (i) applies to arbitrary N-D tensors (images, videos, spectrograms, multivariate time series, tabular tensors), (ii) naturally supports both compression and expansion per mode $(H_{i}\lessgtr D_{i})$, (iii) creates insertion points for normalization/activation between mode maps, and (iv) preserves a clear axis-wise inductive bias without requiring handcrafted groups or convolutional kernels.
Known failure modes and how NdLinear relates.
Axis-factorized operators can underperform when tasks require strong entangled cross-axis interactions (e.g., patterns tied jointly to $(H,W)$ rather than separably to each). Grouped/depthwise convs may also struggle if the chosen grouping misaligns with semantics; axial attention can be sensitive to axis order and depth. NdLinear shares the core trade-off (axis-wise separability vs. full expressivity) but makes it explicit and tunable by per-mode widths/ranks or stacking; it inherits the efficiency benefits of factorization while remaining N-D agnostic in form.
A.3 Tensor Contraction and Regression Layers
Tensor Contraction Layers (TCL).
TCLs [Cichocki, 2014, Novikov et al., 2015] implement a one-shot multilinear map by contracting an input tensor $X\!\in\!\mathbb{R}^{D_{1}\times\cdots\times D_{m}}$ with a set of mode matrices $\{V^{(k)}\in\mathbb{R}^{R_{k}\times D_{k}}\}_{k=1}^{m}$, yielding a reduced tensor $X\times_{1}V^{(1)}\times_{2}\cdots\times_{m}V^{(m)}\in\mathbb{R}^{R_{1}\times\cdots\times R_{m}}$. The goal is typically dimensionality reduction (feature compression) via tensor contractions prior to downstream layers; intermediate positions for normalization/activation and per-mode biases are usually not part of the basic formulation [Cichocki, 2014, Novikov et al., 2015].
Tensor Regression Layers (TRL).
TRLs [Kossaifi et al., 2020] cast prediction as supervised regression/classification with a fixed tensor format for the weight (e.g., Tucker-/TT-structured). Given an input tensor $X$, a TRL fits a low-rank tensor $W$ (plus optional bias) such that $\langle W,X\rangle$ (or a nonlinear variant) matches targets. The emphasis is on learning with low-rank weights for sample/parameter efficiency; the tensor format (ranks/cores) is chosen a priori and does not expose interleaved normalization/activation between modes [Kossaifi et al., 2020].
NdLinear vs. TCL/TRL.
NdLinear is a structure-preserving, learnable linear layer that applies sequential mode-wise maps to activations, $Y\;=\;X\times_{1}W_{1}\times_{2}\cdots\times_{m}W_{m}\in\mathbb{R}^{H_{1}\times\cdots\times H_{m}},$ with an implicit rank-1 Tucker/Kronecker weight and $\sum_{k}D_{k}H_{k}$ parameters. This design differs in intent and mechanics:
  • Goal. TCL primarily contracts modes for reduction [Cichocki, 2014, Novikov et al., 2015]; TRL fixes a low-rank regression format [Kossaifi et al., 2020]; NdLinear is a drop-in linear alternative for N-D activations that preserves the full tensor shape and supports expansion or compression per mode $(H_{k}\gtrless D_{k})$.
  • Execution. TCL performs a single contraction; NdLinear performs sequential per-mode maps, creating natural insertion points for LayerNorm/Dropout/activations and allowing per-mode biases. TRL optimizes a fixed low-rank weight but does not expose interleaved, mode-by-mode transforms during the forward pass.
  • Expressivity/control. All three impose structured priors; NdLinear’s axis-wise separability can be tuned via widths/ranks or stacked blocks, interpolating between strong separability and near-dense behavior while retaining N-D outputs.
TCL as a special case of NdLinear.
Under the joint constraints (i) $H_{k}{=}R_{k}$ for all $k$ (no expansion beyond the contracted size), (ii) no per-mode biases, and (iii) no interleaved operations between mode maps (i.e., a single, commutative product), the one-shot TCL contraction coincides with NdLinear [Cichocki, 2014, Novikov et al., 2015]. Outside this narrow corner, NdLinear is strictly more general and practical: it preserves N-D structure through sequential maps, supports per-mode biases and interleaving (stability), and flexibly expands or compresses each mode.
A.4 Structured Linear Layers
NdLinear sits within the broader family of structured linear layers for parameter efficiency [Denil et al., 2013, Wei et al., 2024]. Classic approaches constrain a large dense weight $W\!\in\!\mathbb{R}^{(\prod_{i}D_{i})\times(\prod_{i}H_{i})}$ by imposing algebraic structure on $W$ itself—e.g., low-rank factorizations [Sainath et al., 2013], block/Butterfly/Monarch-style sparse–fast transforms [Dao et al., 2022], or other learned structured matrices [Sindhwani et al., 2015, Potapczynski et al., 2024]. These methods decouple the parameterization of $W$ from the native organization of the activations, often yielding strong compression but requiring the model to implicitly discover how that structure aligns with the data.
In contrast, NdLinear derives its structure from the N-D layout of the input. Given $X\!\in\!\mathbb{R}^{D_{1}\times\cdots\times D_{n}}$, NdLinear replaces the dense map $\mathbb{R}^{\prod_{i}D_{i}}\!\to\!\mathbb{R}^{\prod_{i}H_{i}}$ with sequential mode-wise transforms $D_{i}\!\to\!H_{i}$, preserving tensor shape and inducing an explicit Kronecker (rank-1 Tucker) weight with only $\sum_{i}D_{i}H_{i}$ parameters. This data-centric factorization (i) makes the axis-wise inductive bias transparent and tunable (via per-mode widths/ranks or stacking), (ii) supports per-mode expansion or compression $(H_{i}\gtrless D_{i})$, and (iii) exposes natural insertion points for normalization/activation between modes—while retaining the efficiency benefits typical of structured linear layers [Denil et al., 2013, Wei et al., 2024, Sainath et al., 2013, Dao et al., 2022, Sindhwani et al., 2015, Potapczynski et al., 2024].
A.5 Graph Neural Networks
Graph Neural Networks (GNNs) address data with irregular connectivity by propagating information over edges via message passing [Scarselli et al., 2008, Micheli, 2009, Bronstein et al., 2017, Zhou et al., 2020, Wu et al., 2020]. Concretely, a GNN updates node features by aggregating messages from neighbors and applying learnable transforms [Gilmer et al., 2017, Kipf and Welling, 2017, Hamilton et al., 2017]. Stacking layers increases the receptive field and enables global interaction, but typically requires multiple rounds of propagation to mix distant nodes [Battaglia et al., 2018].
While an N-D grid (e.g., image or tensor lattice) can be modeled as a graph (one node per cell, edges to local neighbors), this introduces unnecessary overhead on regular grids: message passing is inherently local, so achieving global mixing along each axis often demands many layers, with added memory/compute and potential optimization issues (e.g., depth-related bottlenecks). Moreover, parameter sharing in GNNs is tied to edge types and neighborhood schemas, not directly to axis-wise tensor structure.
NdLinear takes the complementary route for regular tensor grids. Given $X\!\in\!\mathbb{R}^{D_{1}\times\cdots\times D_{n}}$, it applies global, mode-wise linear maps $D_{i}\!\to\!H_{i}$ in a single layer, mixing information along entire axes without constructing a graph or iterating local messages. This preserves the N-D layout, yields $\sum_{i}D_{i}H_{i}$ parameters via an implicit Kronecker (rank-1 Tucker) structure, and exposes insertion points for normalization/activation between mode maps. In short: GNNs excel when connectivity is irregular or non-Euclidean [Bronstein et al., 2017, Zhou et al., 2020], whereas NdLinear specializes in axis-aware, structure-preserving transformations on regular tensors, providing simpler and often faster global mixing along each dimension.
When to use which.
Use GNNs for arbitrary graphs, heterogeneous edge types, and relational reasoning on non-grid data [Scarselli et al., 2008, Micheli, 2009, Bronstein et al., 2017, Zhou et al., 2020, Wu et al., 2020]. Use NdLinear when data are naturally arranged as regular tensors and you want efficient, axis-wise global interactions without flattening; its bias toward axis separability provides parameter savings and predictable behavior on grid-structured domains.
A.6 Other Specialized Approaches
A variety of specialized architectures preserve structure without resorting to full flattening:
Slicing-based layers.
Methods that slice inputs along spatial/temporal (or rotated) subdomains process each slice with shared weights, then recombine [Shao et al., 2016, Dieleman et al., 2016]. This preserves locality and orientation information with modest compute. Limitations: boundaries between slices can hinder cross-slice interaction; designs are task-/geometry-specific and often require bespoke preprocessing.
Capsule Networks.
Capsules use vector/matrix-valued units and routing to model part–whole hierarchies and pose relationships [Hinton et al., 2011, Sabour et al., 2017, Hinton et al., 2018]. They maintain structured representations through learned agreement between capsules. Limitations: routing adds iterative, nontrivial overhead; scaling to large resolutions and datasets has proven challenging; design choices (routing, capsule size) are sensitive.
Hadamard/Fourier feature mixing.
Fixed orthogonal or Fourier-like transforms provide global mixing with $\mathcal{O}(N\log N)$ or even $\mathcal{O}(N)$ cost (e.g., Random Features, Fastfood, FNet, Block-based variants) [Rahimi and Recht, 2007, Le et al., 2013, Tancik et al., 2020, Lee-Thorp et al., 2022, Pan et al., 2022]. Limitations: transforms are fixed (non-learnable) or only weakly parameterized, so alignment with data structure must be recovered by subsequent layers; expressivity depends on depth.
Relation to NdLinear.
NdLinear applies learned, factorized linear maps along each tensor mode, preserving the full N-D layout while enabling axis-wise global mixing with $\sum_{i}D_{i}H_{i}$ parameters. Unlike slicing [Shao et al., 2016, Dieleman et al., 2016], it does not require hand-crafted partitions; unlike capsules [Hinton et al., 2011, Sabour et al., 2017, Hinton et al., 2018], it avoids iterative routing; unlike fixed Hadamard/Fourier mixers [Rahimi and Recht, 2007, Le et al., 2013, Tancik et al., 2020, Lee-Thorp et al., 2022, Pan et al., 2022], it learns mode-wise transforms end-to-end. This yields a simple, geometry-agnostic mechanism for structure-preserving, parameter-efficient linear transformation on regular tensors.
Appendix B More Related Work
Modern neural networks contain substantial parameter redundancy: a large fraction of weights can be predicted from a small subset, sometimes up to 95% with no loss in accuracy [Denil et al., 2013]. This has motivated a broad line of work on efficient parameterizations that preserve accuracy while reducing storage and compute.
Structured tensor factorization.
A major thread leverages high-order structure via tensor decompositions of weights. CP/Tucker-style compressions applied to convolutional kernels reduce parameters and inference cost [Lebedev et al., 2015]. Tensor Train (TT) layers compress fully-connected mappings into compact tensorized operators [Novikov et al., 2015]. Block-Term (BT) tensor networks combine Tucker- and CP-like structure for additional flexibility [Ye et al., 2020]. These tensor-structured layers reduce parameters while retaining rich multi-way interactions by factoring weights across modes.
Structured matrices and parameter sharing.
Another approach imposes algebraic structure on large dense matrices, replacing them with families that admit fast transforms and fewer degrees of freedom. Toeplitz-like and related structured operators provide strong compression with competitive accuracy [Sindhwani et al., 2015]; related families (e.g., circulant, block-circulant) and low-rank factorizations likewise trade unrestricted expressivity for parameter/compute efficiency [Lebedev et al., 2015].
Multi-space representations.
Complementary to structural compression, multi-space learning embeds features into multiple geometries to better capture hierarchy and long-range relations. For example, jointly using Euclidean and hyperbolic spaces for LiDAR yields improved hierarchical encoding and pose estimation [Wang et al., 2023]. Such representations enhance expressivity without necessarily increasing individual layer sizes.
Preserving high-order structure in practice.
Operational layers that respect native tensor axes often strike favorable accuracy–efficiency trade-offs. Depthwise separable convolutions split channel-wise and spatial mixing to cut FLOPs while preserving inductive bias [Chollet, 2017b, Howard et al., 2017]. However, many fully-connected stages still flatten activations, discarding axis structure learned upstream.
Positioning NdLinear.
NdLinear aligns with these trends but differs in where structure is imposed: rather than factorizing weights after flattening, it performs mode-wise learned linear maps directly on N-D activations, preserving tensor layout throughout. Conceptually, it is a rank-1 Tucker (Kronecker) parameterization of the dense linear map, with parameters scaling as $\sum_{i}D_{i}H_{i}$ rather than $\prod_{i}D_{i}\prod_{i}H_{i}$. This data-aligned factorization complements tensorized weights [Lebedev et al., 2015, Novikov et al., 2015, Ye et al., 2020] and structured matrices [Sindhwani et al., 2015], and, like depthwise separable convolutions [Chollet, 2017b, Howard et al., 2017], leverages axis-aware inductive bias—without resorting to flattening.
Appendix C Proofs and Technical Details
C.1 VC-Dimension Analysis
We analyze the expressive capacity of NdLinear compared to standard linear layers. Following Bartlett et al. [2019], any piecewise-linear feedforward network with $P$ parameters has VC-dimension $\Theta(P\log P)$.
Theorem C.1 (VC-Dimension of NdLinear).
Consider input tensors of shape $(B,a,b,c)$ transformed to outputs of shape $(B,d,d,d)$. Let:
  • $N_{\text{vanilla}}=dabc$ (parameters in vanilla linear layer)
  • $N_{\text{nd}}=d(a+b+c)$ (parameters in NdLinear)
Then:
  1. 1.
    As $d\to\infty$ with $a,b,c$ fixed: $N_{\text{nd}}=\Theta(N_{\text{vanilla}})$
  2. 2.
    NdLinear’s VC-dimension is $\Theta(N_{\text{nd}}\log N_{\text{nd}})$
Proof.
For part (1), observe that:
$N_{\text{nd}}=d(a+b+c)=\frac{a+b+c}{abc}\cdot dabc=\frac{a+b+c}{abc}\cdot N_{\text{vanilla}}.$
Since $\frac{a+b+c}{abc}$ is a positive constant (for fixed $a,b,c$), we have:
$N_{\text{nd}}=\Theta(N_{\text{vanilla}})\quad\text{as }d\to\infty.$
For part (2), by the Bartlett et al. result, since NdLinear has $N_{\text{nd}}$ parameters and maintains piecewise-linear structure through ReLU activations:
$\text{VCdim}_{\text{NdLinear}}=\Theta(N_{\text{nd}}\log N_{\text{nd}}).$
Interpretation: While NdLinear uses fewer parameters for finite $d$, as the hidden dimension grows, its parameter count becomes proportional to vanilla linear layers, preserving the same VC-dimension scaling.
Theorem C.2 (Parameter Count Lower Bound).
For positive integers $a,b,c,d$:
$d(a+b+c)>\log(dabc).$
Proof.
We have:
$\log(dabc)$ $=\log d+\log a+\log b+\log c$
$\leq(d-1)+(a-1)+(b-1)+(c-1)\quad\text{(using }\log x\leq x-1\text{)}$
$=d+a+b+c-4.$
Therefore:
$d(a+b+c)-\log(dabc)$ $\geq d(a+b+c)-(d+a+b+c-4)$
$=(d-1)(a+b+c)-d+1+4$
$=(d-1)(a+b+c-1)+4$
$\geq 4>0,$
since $d\geq 1$ and $a+b+c\geq 3$ for non-trivial tensors. ∎
C.2 Peak Activation Memory Analysis
Proposition C.1 (Peak Memory Overhead Bound).
For an input tensor with $m$ modes of sizes $(d_{1},\ldots,d_{m})$ where $\prod_{i=1}^{m}d_{i}=D$, and output sizes $k_{i}\leq d_{i}$, the peak additional activation memory for backpropagation satisfies:
$\frac{\text{extra activation memory}}{\text{baseline activation memory}}\leq\frac{\max_{i}(k_{i}/d_{i})\cdot\min_{i}d_{i}}{m\cdot\min_{i}d_{i}}\leq\frac{1}{m}.$
For typical 3D tensors (m=3), this overhead is at most 33%.
Proof.
During forward pass, NdLinear sequentially transforms each mode. For backpropagation, we store intermediate activations after each mode transformation.
After transforming $j$ modes, the tensor has shape:
$(B,k_{1},\ldots,k_{j},d_{j+1},\ldots,d_{m}).$
The peak extra memory occurs at the stage with the largest intermediate tensor. Since $k_{i}\leq d_{i}$, each intermediate tensor has at most $BD$ elements. The baseline memory is also $BD$ elements.
In the worst case where all $k_{i}=d_{i}$, we have at most (m1)(m-1) intermediate tensors to store, but only one is needed at any given time during backprop (due to sequential processing). Therefore:
$\text{Peak overhead}=\frac{BD\cdot\max_{i}(k_{i}/d_{i})}{BD}\leq\frac{1}{m}.$
C.3 Computational Complexity
Proposition C.2 (Exact FLOP Count).
NdLinear transforming $\mathcal{X}\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$ to $\mathcal{Y}\in\mathbb{R}^{B\times H_{1}\times\cdots\times H_{N}}$ requires:
$\text{FLOPs}_{\text{NdLinear}}=2B\sum_{k=1}^{N}\left[\left(\prod_{j=1}^{k-1}H_{j}\right)\left(\prod_{j=k+1}^{N}D_{j}\right)D_{k}H_{k}\right],$
where the factor of 2 accounts for multiply-add operations.
This is typically $\mathcal{O}(BND^{N+1})$ for $D_{i}=H_{i}=D$, compared to $\mathcal{O}(BD^{2N})$ for vanilla linear layers—yielding orders of magnitude savings as $N$ increases.
Appendix D Implementation Details and Training Protocol
Here we present implementation details and training protocol of NdLinear (Algorithm 1). Implementing NdLinear in practice involves careful attention to efficiency and compatibility with existing deep learning frameworks. We outline key considerations in the following.
Memory Efficiency. Despite handling high-dimensional tensors, NdLinear is memory-efficient due to its factorized parameterization. The forward pass requires allocating intermediate tensors during each mode transformation (after each linear operation, the tensor has one updated dimension). However, these intermediate allocations are of the same order as the input/output size and significantly smaller than the memory required for a gigantic flattened weight matrix. Modern tensor libraries (PyTorch, TensorFlow) facilitate implementing transpose-reshape-multiply steps without excessive data copying. We ensure in-place operations where possible (e.g., using $\mathtt{view}$ in PyTorch). Operations primarily reuse the input buffer for output as each mode is transformed, ensuring modest peak memory usage.
Parameter Initialization. Each weight matrix $W_{i}$ can be initialized using standard strategies for linear layers (Xavier/Glorot [Glorot and Bengio, 2010] or Kaiming [He et al., 2015] initialization based on fan-in and fan-out). Since $W_{i}$ has fan-in $D_{i}$ and fan-out $H_{i}$, the initialization follows ${\rm Uni}(-\sqrt{\frac{6}{D_{i}+H_{i}}},\sqrt{\frac{6}{D_{i}+H_{i}}})$ for Xavier uniform, or analogous formulas for other initializations. This helps maintain stable gradients across modes. One subtle point is that if $n$ is large, each mode’s weight is relatively small, mitigating the risk of extremely large fan-in. We observed no initialization-specific difficulties; indeed, NdLinear’s parameter reduction may help avoid gradient explosion or vanishing issues in deep networks.
Computational Overhead. Factorized operations (multiple transpose and reshape operations with smaller matrix multiplications) are highly optimized in modern BLAS libraries. Practically, runtime is comparable to or faster than fully-connected layers with similar outputs, due to reduced total FLOPs. Python-level overhead is minimal; the algorithm can be implemented in a single forward function looping over modes. For moderate $n$ (up to 4 or 5 dimensions), this loop is short. Explicit loops or unrolling (NdLinear2d, NdLinear3d, etc.) are feasible, but a simple loop suffices. Autograd engines handle tensor operations seamlessly, allowing standard backpropagation. Each $W_{i}$ receives gradients normally from upstream gradients.
Training Protocols. NdLinear layers can be trained end-to-end with standard optimization algorithms (SGD, Adam) just like standard linear layers. Loss functions depend on the task (cross-entropy, MSE, etc.) and are unaffected by NdLinear. However, because NdLinear significantly reduces parameters compared to fully-connected layers, it tends to overfit less, possibly needing less aggressive regularization. Common techniques remain useful: weight decay (L2 regularization) on weights $W_{i}$, and optionally dropout between layers. Dropout can be applied before or after NdLinear; entries in the output tensor $Y$ can be dropped as usual. Specialized regularizers for factorized weights (norm regularization, orthogonality) may help further restrict solution spaces, though not required.
Optimization and Convergence. Practically, each $W_{i}$ is updated based on a portion of the overall error gradient (due to sequential mode transforms). In experiments, all $W_{i}$ matrices learned smoothly with default optimizer settings. If dimensionality varies significantly across modes, gradient clipping or adaptive learning rates per mode may be beneficial. Throughout our numerical investigations, training dynamics are stable overall — NdLinear layers integrated seamlessly into models without requiring special tuning. Standard protocols (learning rate schedules, early stopping criteria, etc.) used for equivalent models with dense layers apply here.
Appendix E Full Experimental Details and Results
We present full experimental details and more results on LoRA fine-tuning in section˜E.1; language-model pretraining (OPT and BERT) in section˜E.1; time-series prediction (RNN and Transformer) in section˜E.2; tabular data in section˜E.3; and vision tasks (CNN, ViT, and DiT) in section˜E.4. The complete experiment code is available at https://github.com/cyclone-trout/ndlinear_neurips.
E.1 Parameter-Efficient Finetuning with LoRA and NdLinear-LoRA
In our study, we utilized state-of-the-art transformer architectures to investigate the impact of targeted modu le adaptation. We selected two base models, Qwen3-1.7B-Base [Yang et al., 2025] and Meta-Llama-3-8B [Dubey et al., 2024], recognized for their robust performance across various tasks. To focus our adaptations, we targeted specific modules within these models, including q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj. This approach allowed us to enhance model capacity and efficiency selectively.
For the adaptation process, we employed Low-Rank Adaptation (LoRA) techniques [Hu et al., 2022], specifically using both NdLinear LoRA and classic LoRA configurations. We explored a range of alpha values (1, 4, and 8) and rank settings (1, 4, and 8) to determine the optimal configuration for our models. This exploration was critical for understanding how different levels of parameter sharing and scaling affect model performance and generalization.
The training process was conducted using the AdamW optimizer, a choice informed by its effectiveness in managing the complexities of transformer models. We set the learning rate to $1\times 10^{-4}$, which provided a suitable balance between convergence speed and training stability. The batch size was set to 1, a decision that facilitated the use of gradient accumulation to optimize GPU memory usage. To ensure the models could handle a wide variety of inputs, we set the maximum sequence length to 512 tokens. The models were trained over 2 epochs, a duration found to be sufficient for achieving significant performance improvements without excessive computational cost. To ensure reproducibility, we used a random seed of 42 across all experiments.
Our models were fine-tuned using two datasets: Math10K and CommonsenseQA. These datasets were chosen for their ability to challenge the models with both mathematical reasoning and commonsense understanding. For evaluation, we employed a diverse set of benchmark datasets, including GSM8K, MultiArith, ARC-C, ARC-E, and BoolQ. This selection allowed us to assess the models’ generalization capabilities across different types of reasoning tasks.
The entire implementation was carried out on a single NVIDIA H100 GPU, using Hugging Face’s AutoModelForCausalLM framework [^1] https://huggingface.co/docs/transformers/en/model_doc/auto, integrated with our custom NdLinear adapter layer. Datasets were tokenized using the default tokenizer for each model, with padding applied to the eos_token. We employed label masking to exclude prompt tokens from loss computation, ensuring that training focused on the relevant portions of the input. Our implementation leveraged PyTorch, along with Hugging Face Transformers, PEFT, and Accelerate, to facilitate efficient model training and adaptation. Evaluation was performed in a zero-shot setting using greedy decoding, which provided a consistent measure of model performance without the variability introduced by sampling methods.
Open Pretrained Transformer (OPT) [Zhang et al., 2022]. For OPT-Small, which originally contains 124M parameters, replacing the standard linear layers with NdLinear reduces the parameter count to 119M. Similarly, for OPT-Mid, the parameter count decreases from 350M to 337M after the substitution.
Table 5: Perplexity comparison for OPT-Small and OPT-Mid models with Linear vs. NdLinear layers.
Linear NdLinear
OPT-Small (Params) 15.970 (124M) 15.755 (119M)
OPT-Mid (Params) 12.926 (350M) 12.565 (337M)
In Table 5, both the OPT-Small and OPT-Mid models achieve lower perplexity scores after replacing standard linear layers with NdLinear layers, despite having fewer parameters. Moreover, the performance improvement becomes more significant as model size increases, with the perplexity gap widening from $0.215$ in OPT-Small to $0.361$ in OPT-Mid. Figure 2 shows that OPT models with NdLinear feedforward layers achieve lower final training and evaluation losses compared to their counterparts using standard linear feedforward layers.
(a) Training Loss during Pre-Training.
(b) Evaluation Loss during Pre-Training.
Figure 2: Training and evaluation loss curves during OPT model pretraining. NdLinear variants consistently achieve lower loss values. x-axis represents the number of training steps.
Zero-shot Tasks. We also evaluate the OPTs’ pretraining on 10 zero-shot NLP tasks:
  • Natural Language Inference Tasks: CB [De Marneffe et al., 2019]
  • Coreference Resolution Tasks: Winogrande [Sakaguchi et al., 2021]
  • Sentence Completion Tasks: COPA [Roemmele et al., 2011], HellaSwag [Zellers et al., 2019]
  • Word Sense Disambiguation Tasks: WiC [Pilehvar and Camacho-Collados, 2018]
  • Question Answering Tasks: ARC-Easy, ARC-Challenge [Clark et al., 2018], OpenBookQA [Mihaylov et al., 2018], BoolQ [Clark et al., 2019]
  • Commonsense Reasoning Tasks: PIQA [Bisk et al., 2020]
During evaluation, we cast all of the above tasks into a multiple-choice format. Namely, the goal is to select the correct completion from a set of candidate options. For each option, we compute the language model (LM) likelihood of the full input consisting of the context concatenated with the candidate completion. To account for differences in the lengths of candidate completions, we compute the average per-token log-likelihood for each option, following [Brown et al., 2020]. The model’s prediction is taken to be the option with the highest per-token likelihood.
Table 6: Perplexity Score and Zero-Shot Performance on OPT Model with and without NdLinear.
OPT-Small OPT-Mid
     Linear NdLinear      Linear NdLinear
Num of Params 124M 119M 350M 337M
Perplexity 15.970 15.755 12.926 12.565
CB 0.32 0.38 0.50 0.52
Winogrande 0.49 0.51 0.50 0.50
COPA 0.58 0.53 0.54 0.56
HellaSwag 0.26 0.26 0.28 0.29
WiC 0.51 0.50 0.50 0.49
ARC-Easy 0.29 0.30 0.29 0.29
ARC-Challenge 0.24 0.25 0.24 0.23
OpenBookQA 0.32 0.35 0.34 0.34
BoolQ 0.49 0.59 0.60 0.56
PIQA 0.53 0.50 0.53 0.53
BERT [Devlin et al., 2019]. We replace the conventional two-layer linear classification head in BERT with an NdLinear layer followed by a classification layer. The NdLinear transforms have hidden dimensions of $(2,2)$. Each model is trained for 200 epochs with a batch size of 32, a hidden layer size of 128, and a learning rate of 0.005.
Table 7: BERT text classification performance on CoLA and SST-2 datasets. NdLinear improves accuracy and ROC AUC with $\approx$85% fewer parameters in the classification head.
Dataset Method Params (Head) Accuracy ROC AUC
CoLA Linear 1,544 0.7790 $\pm$ 0.0143 0.7127 $\pm$ 0.0264
NdLinear 222 0.7906 $\pm$ 0.0142 0.7405 $\pm$ 0.0209
SST-2 Linear 1,544 0.8872 $\pm$ 0.0079 0.8867 $\pm$ 0.0080
NdLinear 222 0.8933 $\pm$ 0.0093 0.8932 $\pm$ 0.0073
E.2 Time Series
Time Series Forecasting. In our experiments using RNNs, we used a sequence length of 24 and a forecast horizon of 12 for all models. The models were trained for 100 epochs using the Adam optimizer with a learning rate of 0.02 and a batch size of 128. The dataset was split into training, validation, and evaluation sets with proportions of 60%, 20%, and 20%, respectively. We set the hidden size to 96, used a single recurrent layer, and applied a dropout rate of 0.3.
For Transformer-based Forecasting tasks, the experiments were conducted using a time series transformer model with a model dimension and hidden dimension both set to 32, a single transformer layer, and a dropout rate of 0.1. The GELU activation function was employed throughout, and the models were trained for 10 epochs with a batch size of 128 and a learning rate of 0.001. All models optimized using Adam and mean squared error as the loss function.
Time Series Classification. We set the hidden size of all RNN layers to 128 and used 3 recurrent layers, with a batch size of 32 and a learning rate of 0.005. Models were trained for 200 epochs using the Adam optimizer and cross-entropy loss.
E.3 Tabular Data
We compare the performance of the Linear and NdLinear models. The Linear Model uses two linear layers for feature extraction, while the NdLinear Model replaces them with NdLinear layers.
Classification. Target labels were one-hot encoded. The Linear Model utilized fully connected layers with input dimension $[11]$ and hidden dimension $[128]$, followed by ReLU and a final linear output layer. The NdLinear Model used custom NdLinear layers with input dimensions $[11,1]$ and hidden dimensions $[11,64]$, also followed by ReLU and a final linear output layer. Both models are trained over 40 epochs with a batch size of 32 and a learning rate of $0.0001$ using AdamW optimizer. Data was randomly shuffled and split into 80% training and 20% testing sets. Cross-entropy loss is used for training, and classification accuracy is used for model evaluation.
Regression. Target labels were kept as continuous values. The Linear Model utilized fully connected layers with input dimension $[14]$ and hidden dimension $[128]$, followed by ReLU and a final linear output layer. The NdLinear Model used custom NdLinear layers with input dimensions $[14,1]$ and hidden dimensions $[32,64]$, also followed by ReLU and a final linear output layer. Both models are trained over 40 epochs with a batch size of 32 and a learning rate of $0.0002$ using AdamW optimizer. Data was randomly shuffled and split into 80% training and 20% testing sets. MSE loss is used for both training and model evaluation.
Table 8: NdLinear with MLPs on tabular datasets. For classification (Cardio Disease), the metric is Accuracy (higher is better). For regression (Delivery Time), the metric is MSE (lower is better).
Dataset Task Method #Params Metric
Cardio Disease Classif. (Accuracy) Linear 18 306 0.7265
NdLinear 5 962 0.7321
Delivery Time Regress. (MSE) Linear 18 561 70.508
NdLinear 7 873 67.824
E.4 Vision
Image Classification with CNN. The NdLinear version uses three transforms with hidden dimensions of $(64,8,8)$, while the Linear version uses a single hidden dimension of $256$. Models were trained for 50 epochs using Adam optimizer (learning rate $0.001$), batch size 64, cross-entropy loss, and mixed-precision training on CUDA when available.
Table 9: Image classification with CNNs on CIFAR-10 (top-1 Acc.) and CIFAR-100 (top-5 Acc.). NdLinear achieves higher accuracy with fewer parameters.
Dataset #Params Method Accuracy
CIFAR-10 1.07M Linear 0.7426 $\pm$ 0.0025
65k NdLinear 0.7689 $\pm$ 0.0060
CIFAR-100 1.09M Linear 0.6587 $\pm$ 0.0075
433k NdLinear 0.7096 $\pm$ 0.0121
Vision Transformers (ViT) [Dosovitskiy et al., 2021]. Training used a batch size of 512, AdamW optimizer, learning rate $2.75\times 10^{-4}$ for 30 epochs, and a distillation temperature of 3. Input images ($224\times 224$) were augmented with random cropping and horizontal flipping.
Figure 3: NdLinear’s efficiency. Reduced ViT model parameter counts on CIFAR-10 and CIFAR-100 for a distillation task.
Table 10: NdViT vs. Naive ViT: Accuracy and parameter efficiency on CIFAR-10 (Acc@1) and CIFAR-100 (Acc@5). NdViT shows improved accuracy with fewer parameters.
Dataset
Num. of
Transformers
NdViT (Ours) Naive
200 300 400 500
CIFAR10
3 Blocks 65.77 $\pm$ 0.47 67.53 $\pm$ 0.70 69.00 $\pm$ 1.27 62.09 $\pm$ 0.40
6 Blocks 68.48 $\pm$ 0.75 70.20 $\pm$ 0.73 72.03 $\pm$ 0.46 65.19 $\pm$ 0.64
9 Blocks 70.27 $\pm$ 0.35 71.50 $\pm$ 0.58 72.53 $\pm$ 0.54 68.52 $\pm$ 1.24
CIFAR100
3 Blocks 70.78 $\pm$ 1.36 73.10 $\pm$ 1.06 74.14 $\pm$ 1.66 69.34 $\pm$ 0.88
6 Blocks 73.60 $\pm$ 0.83 75.07 $\pm$ 0.14 76.37 $\pm$ 0.71 73.84 $\pm$ 0.39
9 Blocks 74.24 $\pm$ 0.32 75.52 $\pm$ 0.73 76.61 $\pm$ 0.26 75.60 $\pm$ 0.70
Diffusion Transformers (DiT) [Peebles and Xie, 2023]. We used a learning rate of $1\times 10^{-4}$ for $256\times 256$ images, mixed-precision (bfloat16) training, automatic gradient accumulation, and a batch size of 256, varying model depth and attention heads.
Table 11: FID-10k scores for DiT models: Pre-trained vs. NdLinear variants with fewer parameters.
NdLinear NdLinear Baseline
Parameter Count 619M 563M 674M
FID-10k 5.4876 5.9420 5.4109
Figure 4: DiT achieving lower (better) FID scores for image generation on ImageNet-100 when trained from scratch with comparable parameters.
Appendix F Detailed Ablation Studies
F.1 Per-Mode Bias Impact
We ablate per-mode bias terms across widths on the Radius Bump task.
We generate 1,0001{,}000 train and $200$ test samples. Hidden widths {16, 32, 64, 128}. We have NdLinear-MLPs replacing Linear blocks with NdLinear, using per-axis hidden shapes {(4,4), (8,8), (16,16), (32,32)} matched to the hidden widths above. For all models, training uses Adam (learning rate $10^{-3}$) for 4,0004{,}000 epochs under identical schedules, with early stopping when training loss $10^{-4}$; loss is mean squared error (MSE). We report parameter count and test MSE (mean $\pm$ std).
The benefit of per-mode bias grows with width, reaching +15.2% MSE improvement at width 128.
Table 12: Per-mode bias ablation on Radius Bump
Width Bias MSE (mean) MSE (std) Params
16 False 0.00332264 0.00173532 36
16 True 0.00335201 0.00157289 46
32 False 0.00332793 0.00170700 72
32 True 0.00318228 0.00163146 90
64 False 0.00336938 0.00160766 144
64 True 0.00305152 0.00145630 178
128 False 0.00339010 0.00173152 288
128 True 0.00287480 0.00139560 354
F.2 Axis Ordering Sensitivity
We permute axes (original, reverse, random) and measure CIFAR-100 performance retention. Finding: robustness to ordering ($\leq$ 4 pp spread).
Table 13: Axis ordering sensitivity (CIFAR-100)
Variant (axes) Accuracy vs. baseline
Original order 100%
Reverse 99% $\pm$ 1%
Random 96% $\pm$ 1%
F.3 Radius Bump: Last-Layer Entropy, Performance, and Compression
We evaluate Dense MLP vs. NdLinear on the Radius Bump task across three difficulty levels (shell thickness $\sigma\in\{0.10,0.20,0.30\}$). Inputs are $x\in[-1,1]^{10}$ with i.i.d. coordinates $x_{i}\sim\mathrm{U}(-1,1)$; the target is $y=\exp\!\bigl(-(\|x\|-0.8)^{2}/(2\sigma^{2})\bigr).$ For each $\sigma$ we generate 1,0001{,}000 train and $200$ test samples. Architectures: (i) Dense MLP baselines with hidden widths {16, 32, 64, 128} and depths {2, 3}; (ii) NdLinear-MLPs replacing Linear blocks with NdLinear, using per-axis hidden shapes {(4,4), (8,8), (16,16), (32,32)} matched to the dense widths above and depths {2, 3}. For all models, training uses Adam (learning rate $10^{-3}$) for 4,0004{,}000 epochs under identical schedules, with early stopping when training loss $10^{-4}$; loss is mean squared error (MSE). We report parameter count, forward-pass FLOPs (from analytical op counts), test MSE (mean $\pm$ std), and last-layer average output entropy (AveEntropy).
Table 14: Radius Bump ($\sigma=0.10$; thin / hardest). Lower AveEntropy indicates more compact last-layer representations.
Width Kind Depth Hidden Shape Params Train MSE Test MSE $\pm$ Std FLOPs ($\cdot 10^{9}$) AveEntropy
16 dense 2 (16) 193 1.17e-5 0.00086 ± 0.00176 2.78 0.476
16 dense 3 (16) 465 1.37e-6 0.00087 ± 0.00189 6.70 0.625
16 nd 2 (4, 4) 46 1.19e-5 0.00086 ± 0.00186 0.66 0.405
16 nd 3 (4, 4) 86 1.18e-5 0.00086 ± 0.00186 1.24 0.518
32 dense 2 (32) 385 1.18e-5 0.00086 ± 0.00169 5.54 0.459
32 dense 3 (32) 1,441 2.93e-6 0.00084 ± 0.00188 20.75 0.560
32 nd 2 (8, 8) 90 1.20e-5 0.00086 ± 0.00182 1.30 0.321
32 nd 3 (8, 8) 234 1.16e-5 0.00086 ± 0.00178 3.37 0.426
64 dense 2 (64) 769 1.19e-5 0.00086 ± 0.00169 11.07 0.407
64 dense 3 (64) 4,929 1.59e-8 0.00085 ± 0.00196 70.98 0.507
64 nd 2 (16, 16) 178 1.12e-5 0.00086 ± 0.00175 2.56 0.292
64 nd 3 (16, 16) 722 1.12e-5 0.00086 ± 0.00180 10.40 0.377
128 dense 2 (128) 1,537 5.79e-6 0.00089 ± 0.00253 22.13 0.363
128 dense 3 (128) 18,049 1.84e-9 0.00089 ± 0.00158 259.91 0.469
128 nd 2 (32, 32) 354 1.12e-5 0.00086 ± 0.00158 5.10 0.295
128 nd 3 (32, 32) 2,466 8.71e-6 0.00086 ± 0.00204 35.51 0.286
Table 15: Radius Bump ($\sigma=0.20$; medium). NdLinear often attains both lower error and lower last-layer entropy.
Width Kind Depth Hidden Shape Params Train MSE Test MSE $\pm$ Std FLOPs ($\cdot 10^{9}$) AveEntropy
16 dense 2 (16) 193 7.75e-4 0.00256 ± 0.00235 2.78 0.493
16 dense 3 (16) 465 1.29e-5 0.00166 ± 0.00118 6.70 0.592
16 nd 2 (4, 4) 46 9.91e-4 0.00273 ± 0.00273 0.66 0.410
16 nd 3 (4, 4) 86 1.01e-3 0.00275 ± 0.00272 1.24 0.521
32 dense 2 (32) 385 6.03e-4 0.00266 ± 0.00234 5.54 0.461
32 dense 3 (32) 1,441 5.92e-6 0.00180 ± 0.00117 20.75 0.560
32 nd 2 (8, 8) 90 9.92e-4 0.00273 ± 0.00232 1.30 0.323
32 nd 3 (8, 8) 234 1.19e-4 0.00192 ± 0.00089 3.37 0.428
64 dense 2 (64) 769 3.17e-4 0.00368 ± 0.00344 11.07 0.403
64 dense 3 (64) 4,929 1.84e-8 0.00128 ± 0.00190 70.98 0.518
64 nd 2 (16, 16) 178 8.35e-4 0.00252 ± 0.00216 2.56 0.293
64 nd 3 (16, 16) 722 9.56e-6 0.00071 ± 0.00019 10.40 0.344
128 dense 2 (128) 1,537 2.53e-5 0.00454 ± 0.00330 22.13 0.364
128 dense 3 (128) 18,049 2.01e-10 0.00159 ± 0.00102 259.91 0.471
128 nd 2 (32, 32) 354 6.07e-4 0.00233 ± 0.00211 5.10 0.296
128 nd 3 (32, 32) 2,466 2.33e-6 0.00081 ± 0.00019 35.51 0.284
Table 16: Radius Bump ($\sigma=0.30$; thick / easiest). NdLinear depth-3 variants achieve the lowest errors with the lowest entropies.
Width Kind Depth Hidden Shape Params Train MSE Test MSE $\pm$ Std FLOPs ($\cdot 10^{9}$) AveEntropy
16 dense 2 (16) 193 3.65e-3 0.00716 ± 0.00157 2.78 0.505
16 dense 3 (16) 465 1.87e-4 0.00210 ± 0.00084 6.70 0.627
16 nd 2 (4, 4) 46 5.91e-3 0.00827 ± 0.00261 0.66 0.406
16 nd 3 (4, 4) 86 2.73e-3 0.00710 ± 0.00345 1.24 0.527
32 dense 2 (32) 385 2.03e-3 0.00636 ± 0.00209 5.54 0.461
32 dense 3 (32) 1,441 1.42e-5 0.00166 ± 0.00063 20.75 0.588
32 nd 2 (8, 8) 90 3.53e-3 0.00564 ± 0.00190 1.30 0.323
32 nd 3 (8, 8) 234 6.78e-5 0.00045 ± 0.00021 3.37 0.424
64 dense 2 (64) 769 8.76e-4 0.00968 ± 0.00175 11.07 0.406
64 dense 3 (64) 4,929 1.34e-5 0.00152 ± 0.00029 70.98 0.521
64 nd 2 (16, 16) 178 3.01e-3 0.00502 ± 0.00142 2.56 0.293
64 nd 3 (16, 16) 722 9.46e-6 0.00029 ± 0.00021 10.40 0.341
128 dense 2 (128) 1,537 1.66e-4 0.01057 ± 0.00324 22.13 0.362
128 dense 3 (128) 18,049 2.03e-10 0.00249 ± 0.00051 259.91 0.474
128 nd 2 (32, 32) 354 1.78e-3 0.00460 ± 0.00096 5.10 0.296
128 nd 3 (32, 32) 2,466 3.41e-5 0.00041 ± 0.00014 35.51 0.278
Figure 5: $\alpha=0.1$ (hard): Narrow bump, very challenging.
Figure 6: $\alpha=0.2$ (medium): Moderate difficulty.
Figure 7: $\alpha$ = 0.3 (easy): Wide bump, easier to approximate.
Across all $\sigma$, NdLinear variants consistently exhibit lower AveEntropy than parameter-matched dense models at similar widths/depths, while using markedly fewer parameters and FLOPs. At $\sigma=0.20$ and $\sigma=0.30$, depth-3 NdLinear achieves the lowest test errors (e.g., $0.00071\pm 0.00019$ at width 64, and $0.00029\pm 0.00021$ at width 64) together with lower entropies (0.344 and 0.341, respectively) than dense counterparts (0.00128 $\pm$ 0.00190 with entropy 0.518; 0.00152 $\pm$ 0.00029 with entropy 0.521). For the hardest setting ($\sigma=0.10$), NdLinear matches dense test error while maintaining lower entropy and large reductions in parameters and FLOPs. These patterns support the claim that lower last-layer entropy paired with equal-or-better error corresponds to better compression.
F.4 Hyperparameter Sensitivity
Grid search over learning rate, hidden size, and batch size on CIFAR-100. Finding: stable efficiency metrics and competitive accuracy across settings.
Table 17: Hyperparameter sensitivity sweep (CIFAR-100)
LR Hidden Size Batch GFLOPs Latency (s) Params Acc@5 (%)
0.001 128,4,4 128 0.848 0.001415 232,876 72.75
0.001 256,2,2 64 0.901 0.001451 138,760 74.30
0.001 128,4,4 64 0.848 0.001381 232,876 72.67
0.001 128,4,4 32 0.848 0.001409 232,876 72.48
0.001 256,2,2 32 0.901 0.001388 138,760 73.80
0.001 256,2,2 128 0.901 0.001400 138,760 74.75
0.01 128,4,4 128 0.848 0.001318 232,876 71.62
0.01 128,4,4 64 0.848 0.001376 232,876 66.73
0.01 128,4,4 32 0.848 0.001364 232,876 66.56
0.01 256,2,2 64 0.901 0.001257 138,760 64.86
0.01 256,2,2 128 0.901 0.001312 138,760 70.20
0.01 256,2,2 32 0.901 0.001323 138,760 63.75
Note: GPU memory usage constant at 34.04-34.40 MB across all configurations.
F.5 Sample Efficiency
Fix task structure and vary data size; measure samples needed to reach target error at different entanglement levels.
  • Nearly-separable tasks ($\alpha=0.1$): NdLinear reached target MSE with only 2,000 samples; parameter-matched linear model required over 10,000 samples.
  • Highly-entangled tasks ($\alpha=0.9$): Standard linear model achieved target MSE with 15,000 samples; NdLinear struggled to match this performance even with 25,000 samples.
F.6 Training and Memory Overhead
Measure peak activation memory and per-epoch training time after replacing a single dense GEMM with mode-wise GEMMs. Finding: empirical overheads are small ($3\%$ memory, $2\%$ time) as seen in Table 18.
Table 18: Empirical overheads across architectures
Model Peak Mem (MB) Epoch Time (s)
CIFAR-100 CNN 35.17 $\rightarrow$ 36.91 (+2.0%) 47.2 $\rightarrow$ 47.8 (+0.6%)
ETTh1 RNN 32.58 $\rightarrow$ 33.41 (+1.2%) 12.3 $\rightarrow$ 12.6 (+1.2%)
Vision Transformer 127.3 $\rightarrow$ 130.1 (+1.1%) 179 $\rightarrow$ 185 (+1.6%)
F.7 Comparison with Alternative Structured Layers
CNN head on CIFAR-100 comparing NdLinear to TRL/TCL and TT. Finding: NdLinear achieves higher Acc@5 with fewer params, lower FLOPs, and lower latency.
Table 19: NdLinear vs. TRL/TCL vs. $T$ (CIFAR-100 Acc@5)
Method Mem (MB) Acc@5 Latency (s) FLOPs (G) Params
NdLinear 35.16 0.7133 0.000976 0.843 433,588
TRL/TCL 35.60 0.6935 0.001116 3.97 548,032
$T$ 100.44 0.5617 0.005871 5.25 769,316